### v2

# Concept-1K: A Novel Benchmark for Instance Incremental Learning

Abstract
Incremental learning (IL) is essential to realize the human-level intelligence in the neural network. However, existing IL scenarios and datasets are unqualified for assessing forgetting in PLMs, giving an illusion that PLMs do not suffer from catastrophic forgetting. To this end, we propose a challenging IL scenario called instance-incremental learning (IIL) and a novel dataset called Concept-1K, which supports an order of magnitude larger IL steps. Based on the experiments on Concept-1K, we reveal that billion-parameter PLMs still suffer from catastrophic forgetting, and the forgetting is affected by both model scale, pretraining, and buffer size. Furthermore, existing IL methods and a popular finetuning technique, LoRA, fail to achieve satisfactory performance. Our study provides a novel scenario for future studies to explore the catastrophic forgetting of PLMs and encourage more powerful techniques to be designed for alleviating the forgetting in PLMs. The data, code and scripts are publicly available  [^1] https://github.com/zzz47zzz/codebase-for-incremental-learning-with-llm.
Concept-1K: A Novel Benchmark for Instance Incremental Learning
Junhao Zheng, Shengjie Qiu, Qianli Ma* School of Computer Science and Engineering, South China University of Technology, Guangzhou, China junhaozheng47@outlook.com, shengjieqiu6@gmail.com, qianlima@scut.edu.cnthanks: *Corresponding author
1 Introduction
Incremental Learning (IL) aims to equip models with new knowledge without forgetting the old knowledge. It is of critical importance for research on the verge of human-level intelligence. Recent years have witnessed significant breakthroughs of Pretrained Language Models OpenAI (2023) (PLMs) in not only downstream tasks in natural language processing but also in real-world applications. Although there is a wealth of research on incremental learning with PLMs, we still lack a fundamental understanding of the incremental learning ability of PLMs. For instance, we can not clearly answer the questions “Do PLMs still suffer from catastrophic forgetting?”, “Do larger PLMs forget less than the smaller ones?”, “What is the role of pretraining in incremental learning?”. The root cause is the lack of suitable IL benchmarks designed for PLMs.
Figure 1: The illustration of Instance Incremental Learning (IIL).
There are primarily three IL scenarios for PLMs: class-incremental learning (CIL), task-incremental learning (TIL), and continual pretraining (CPT). However, none of them are appropriate to evaluate the IL ability of PLMs. CIL is designed for classification tasks such as text classification, and its goal is to learn new classes incrementally. On the one hand, state-of-the-art (SOTA) PLMs with billion-level parameters are overqualified for the above classification tasks with only dozens of categories. On the other hand, the pretraining corpus is likely to contain the knowledge required for the classification tasks (data leakage issue). As shown empirically by Zheng et al. (2023), sequential training frozen PLMs with expanding classifiers yields comparable or even superior performance with SOTA IL methods. TIL aims to learn new tasks incrementally Sun et al. (2020); Qin and Joty (2022). Apart from the data leakage issue, the diversity of tasks and orders across research makes it difficult to readily and fairly compare IL algorithms. The last scenario CPT aims at continual pretraining models on the corpus from different domains. However, the evaluation relies on the performance of downstream tasks, where we can hardly identify what knowledge is learned or forgotten.
In this paper, we propose a novel incremental learning scenario called Instance Incremental Learning (IIL), for benchmarking the IL ability of SOTA PLMs. Unlike the IL scenario mentioned above, IIL regards each concept as an instance and is more practical and challenging for existing PLMs. Specifically, we are motivated by the human learning process and expect PLMs to learn new concepts incrementally without forgetting. For example, in Figure 1, humans can learn new concepts that are constantly emerging, such as “Metaverse” and “Quantum Computing”. After learning more concepts such as “Web3.0” and “Non-Fungible Token”, humans will not immediately forget the previously learned concepts such as “Metaverse”. In contrast, we will later empirically show that PLMs with billions of parameters fail to recall the knowledge of old concepts.
Additionally, we present a dataset called Concept-1K for the IIL scenario. Concretely, we utilize the vast knowledge of GPT4 OpenAI (2023) to generate recently emerged concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical, resulting in a total of 1023 concepts. Next, we prompted GPT4 to construct triplets for each concept with the relation in ConceptNet Speer et al. (2017) where each triplet instantiates a knowledge point. Finally, we obtain 16653 training-test QA pairs corresponding to 16653 knowledge points from 1023 novel concepts.
We conduct experiments on different scales of PLMs with various pretraining steps, buffer size, and finetuning methods under the IIL scenario. The results show that PLMs suffer from catastrophic forgetting seriously and have a huge gap to human beings. In addition, we reveal that existing rehearsal-free IL methods fail to achieve superior performance. Furthermore, we find that in-context learning fails to incrementally adapt PLMs to novel concepts. In summary, the contribution is two-fold:
  • We propose a novel incremental learning scenario called IIL, and a dataset called Concept-1K for better evaluating the IL ability of PLMs.
  • Based the experiments on Concept-1K, we reveal that PLMs still suffer from catastrophic forgetting and existing techniques fail to achieve superior performance when no old data is stored.
2 Instance Incremental Learning
Problem Formulation. Formally, IIL aims to learn a model $f_{\theta}:\mathbf{x}\rightarrow\mathbf{y}$ from the sequence of concepts $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n},\cdots,\mathcal{C}_{N}\}$, where $N$ is the number of concepts, and both input $\mathbf{x}$ and output $\mathbf{y}$ are natural language. The $n$-th concept $\mathcal{C}_{n}$ contains $M_{n}$ training-test pairs $\mathcal{D}^{(n)}=\{\mathbf{x}_{i}^{(n),train},\mathbf{x}_{i}^{(n),test},\mathbf{y}_{i}^{(n)}\}_{i=1}^{M_{n}}$, where $\mathbf{x}_{i}^{(n),train}$ and $\mathbf{x}_{i}^{(n),test}$ are the training and test inputs, and $\mathbf{y}_{i}^{(n)}$ is the target output of them. Each training-test pair corresponds to the same knowledge point about the concept $\mathcal{C}_{n}$. For practical training and evaluation, we evenly divide $N$ concepts into $T$ ($T\leq N$) groups corresponding to $T$ incremental steps. The model will be evaluated after learning each group of concepts.
Metric. We adopt four evaluation metrics for IIL, including Memorization Accuracy (MA), Memorization Forgetting rate (MF), Generalization Accuracy (GA), and Generalization Forgetting rate (GF). Specifically, MA and MF measure how much knowledge of training samples is memorized and forgotten. GA and GF measure how much knowledge is generalized to the test samples and is forgotten. The formal definition is provided in Appendix A.
Table 1: Comparison between Concept-1K and the widely-used datasets for incremental learning with PLMs. Concept-1k supports an order of magnitude larger incremental learning steps than existing ones.
Dataset # Classes / Concepts Task Type
AGNews Zhang et al. (2015) 4 Topic Classification
DBPedia Zhang et al. (2015) 14
YaHoo Zhang et al. (2015) 10
CLINC150 Larson et al. (2019) 150 Intent Classification
Banking77 Casanueva et al. (2020) 77
FewRel Han et al. (2018) 80 Relation Extraction
TACRED Zhang et al. (2017) 40
Few-NERD Ding et al. (2021) 66 Named Entity Recognition
Ontonotes5 Hovy et al. (2006) 18
I2B2 Murphy et al. (2010) 16
Concept-1K 1023 Question and Answering
(a) Scale vs Buffer
[Pretraining Step=Final]
(b) Pretraining vs Scale
[Buffer Size=0]
(c) Pretraining vs Scale
[Buffer Size=2000]
(d) Pretraining vs Scale
[Buffer Size=ALL]
(e) Scale vs Buffer
[Pretraining Step=Final]
(f) Pretraining vs Scale
[Buffer Size=0]
(g) Pretraining vs Scale
[Buffer Size=2000]
(h) Pretraining vs Scale
[Buffer Size=ALL]
Figure 2: The analysis of memorization (top row) and generalization (bottom row) accuracy on Concept-1K. The backbone model is in {Pythia-70M, 160M, 410M, 1B, 1.4B, 2.8B}. The pretraining step is in {0, 16, 128, 1000, 10000, 143000 (final version)}. Each point represents the result of sequential finetuning under the IIL scenario. The detailed results with standard deveriation are provided in Table 7, 8, 9, 10, 11, 12, 13, and 14.
Concept-1K To avoid the data leakage issue, we collect novel concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical. To begin with, we generate 600 concepts for each domain with GPT4. We then manually filter out outdated, vague, or imaginary concepts and select the latest, most specific, and most informative concepts, resulting in a total of 1023 concepts. The concept list is provided in Table 17. Next, we use triplets to represent “knowledge” and prompted GPT4 to construct 20 triplets for each concept with the relation in ConceptNet Speer et al. (2017). To avoid conflict, we filter out the triplets with the same concept and relation. In addition, we filter out the triplets with relations such as “RelatedTo” and “HasContext” to ensure specificity. Finally, we utilize GPT4 to convert each triplet into a pair of training and test instances in QA format. We provide the examples and statistics of Concept-1K in Table 1, 15, and Figure 5, 6.
3 Experimental Results
We divide 1023 concepts of Concept-1K into 10 groups for IIL. We explore the impact of the model scale, pretraining, buffer size, and finetuning methods (FULL or LoRA) on the performance under the IIL scenario. We also investigate whether in-context learning is qualified to adapt PLMs to novel concepts in Concept-1K. In addition, we analyze the IIL performance at the concept level in the Appendix C.4. The detailed experimental settings, implementation details, and introduction of baselines are provided in the Appendix B.
Table 2: The forgetting of PLMs with different scales. The pretraining step is final and buffer size is 0.
160M 410M 1B 1.4B 2.8B
MF ($\downarrow$) 76.07±1.08 65.19±0.31 55.36±0.21 56.70±0.24 51.12±0.42
GF ($\downarrow$) 5.35±0.59 15.39±0.16 18.22±0.07 21.70±0.07 23.97±0.18
The Role of Model Scale and Buffer Size. The results are summarized in Figure 2 and Table 2. The findings are summarized as follows: (1) Figure 2 (a) and (e) show that a larger buffer size or a larger PLM improve the accuracy of memorization and generalization. (2) The model scale determines the upper limit of the generalization performance. (3) However, the memorization accuracy of the 2.8B model is still unsatisfactory when no buffer is used. It indicates that the billion-parameter PLMs still suffer from catastrophic forgetting, and data replay is an effective technique for alleviating it. (4) Table 2 shows that when PLMs become larger, the memorization forgetting decreases while the generalization forgetting increases. It indicates that larger PLMs forget less training samples while forget more generalized knowledge since more knowledge are generalized.
The Role of Pretraining. Figure 2 (b) shows that the memorization performance increases at the early stage of pretraining (step 0 - step 10000). It indicates that pretraining enhances the memorization ability of PLMs to novel concepts. However, when performing more pretraining steps, the memorization performance will be degraded. In contrast, Figure 2 (f) shows that the generalization performance increases monotonically for PLMs larger than 160M. The reason may be that PLMs will gradually learn to extract the knowledge behind the text during the pretraining process instead of simply remembering specific texts. In addition, especially for larger models, pretraining is more beneficial to generalization ability rather than memorization ability. Furthermore, Figure 2 (c)(d)(g)(f) indicates that a larger buffer size improves both memorization and generalization ability for all pretraining steps. The improvement of memorization ability is more significant than that of generalization ability.
Table 3: The performance of SOTA methods on Concept-1K. The detailed results are in Figure 4.
Method MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
SEQ 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
EWC Kirkpatrick et al. (2017) 59.83±0.62 18.09±0.28 62.46±0.69 15.07±0.33
LAMOL_g Sun et al. (2020) 58.76±0.53 15.35±0.46 64.64±0.59 13.61±0.47
LAMOL_t Sun et al. (2020) 58.29±2.17 15.24±0.37 66.66±2.42 14.05±0.38
L2KD Chuang et al. (2020) 28.34±0.29 10.87±0.01 32.45±0.63 8.55±0.42
PCLL Zhao et al. (2022) 61.94±1.41 20.06±0.42 63.04±1.57 16.27±0.37
LFPT5 Qin and Joty (2022) 0.63±0.04 0.84±0.01 0.04±0.03 0.03±0.01
LAMOL_KD Zheng et al. (2023) 72.33±0.45 18.20±0.37 49.25±0.32 10.61±0.42
REPLAY (buffer size=2000) 77.31±0.22 22.48±0.29 46.97±1.50 10.57±0.24
REPLAY (buffer size=Alll) 99.01±0.11 25.70±0.44 0.70±0.16 1.44±0.88
Performance of SOTA Methods. We compare 8 rehearsal-free methods and data replay (REPLAY) under the IIL scenario. Table 3 indicates that although existing methods have improved sequential finetuning (SEQ), there is still a gap in the data replay with only 2000 samples (about 12 % of the total sample). The gap in memorization accuracy is much greater than generalization accuracy, indicating that PLMS still suffers from catastrophic forgetting under the challenging IIL scenario. The result urges us to design a more powerful IL algorithm to reduce the need for data replay.
Table 4: The IIL performance of full finetuning (FULL) and LoRA on various backbones.
MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
Pythia-410M (FULL) 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
Pythia-410M (LoRA) 15.72±0.72 6.58±0.15 30.97±1.15 2.96±0.14
Pythia-2.8B (LoRA) 36.56±0.93 10.93±0.17 83.94±0.64 6.04±0.23
Vicuna 7B (LoRA) 42.28±0.25 16.74±0.33 78.32±0.54 7.69±0.47
LLaMa 7B (LoRA) 41.76±0.27 16.20±0.18 80.55±0.39 8.08±0.31
LLaMa 13B (LoRA) 48.90±0.68 22.67±0.30 74.55±0.46 12.26±0.18
LoRA v.s. Full Finetuning. LoRA Hu et al. (2021) only finetunes a small proportion of parameters (usually less than 1%) and effectively adapt PLMs to specific domains. We compare the performance of full finetuning with LoRA in the IIL setting. Table 4 shows that LoRA limits the ability to memorize new knowledge or generalize to new instances. More specifically, the memorization and generalization accuracy of Pythia-410M is higher than LLaMa 7B (LoRA). This reminds us that when we need PLM to learn much new knowledge, we should prioritise using full fine-tuning instead of LoRA.
Table 5: The accuracy of in-context learning on the full Concept-1K (not IIL). “Rand.”, “Same Conc.”, and “Same Know.” represent that the demonstration samples are selected randomly, or from the instances related to the same concept, or from the instance related to the same knowledge (i.e., the same training-test pair).
1 Shot 5 Shot
Rand. Same Conc. Same Know. Rand. Same Conc. Same Know.
Pythia-70M 0.02±0.00 0.10±0.04 91.78±0.00 0.05±0.02 0.17±0.01 15.02±0.00
Pythia-160M 0.10±0.02 0.15±0.02 41.36±0.00 0.37±0.02 0.55±0.15 13.49±0.00
Pythia-410M 0.49±0.17 0.75±0.23 40.66±0.00 2.32±0.80 2.04±0.63 18.45±0.00
Pythia-1B 0.88±0.36 1.21±0.13 43.73±0.00 3.03±0.46 2.69±0.56 35.75±0.00
Pythia-1.4B 1.92±0.28 2.52±0.16 54.37±0.00 4.56±0.58 3.80±0.01 45.59±0.00
Pythia-2.8B 1.81±0.45 2.50±0.27 53.56±0.00 5.69±0.35 4.07±0.93 60.20±0.00
LLaMa 7B 4.32±0.11 4.93±0.72 83.57±0.00 8.79±0.16 6.24±0.23 66.25±0.00
Vicuna 7B 6.67±0.74 7.00±0.88 55.67±0.00 9.63±0.78 7.75±0.15 36.04±0.00
GPT 3.5 6.60 8.20 51.60 8.60 13.00 74.80
GPT 4 10.20 10.40 76.60 7.40 21.80 86.20
In-Context Learning. Due to computing resources and privacy concerns, it is impractical to fully or partially fine-tune SOTA PLMs such as GPT4. We evaluate the in-context learning performance on the whole Concept-1K to investigate the ability to learn novel concepts. Table 5 indicates that the performance is unsatisfactory for all PLMs when the demonstration instances are less related to the test instance. Additionally, we can improve the performance by leveraging demonstration instances that share the same knowledge as the test instances. In other words, PLMs achieves superior performance only when the demonstration instances containing exactly the same knowledge as the test sample. Therefore, in-context learning can not meet the goal of adapting PLMs to new knowledge.
4 Conclusion
Incremental learning is a research field that has developed rapidly in recent years. However, the existing IL scenario and datasets are not suitable for SOTA PLMs and even give an illusion that PLMs do not suffer from catastrophic forgetting. This study presents a novel IL scenario called instance-incremental learning and a novel dataset called Concept-1K, supporting an order of magnitude larger steps for IL. We hope our study will further inspire follow-up works to explore the catastrophic forgetting in PLMs and to design more powerful techniques for finetuning and incremental learning.
Limitations
There are two limitations of this research: (1) The knowledge of Concept-1K is defined in the form of triplets, which can not cover the knowledge in a broad sense. (2) Apart from the experiments of in-context learning, other experiments are conducted on PLMs with less than 13B parameters. The conclusion of these experiments may not hold when finetuning SOTA PLMs such as GPT4 with more than 100B parameters.
Ethical Considerations
The ethical considerations of our research are carefully addressed to ensure compliance with relevant standards and transparency. To this end, we provide the following clarifications:
1. Dataset Collection: Our research employs GPT4 to construct Concept-1K and filter out offensive or harmful instances. The use of GPT4 was consistent with their intended use. The dataset Concept-1K is publicly available for academic and research purposes.
2. Reproducibility: We provide a detailed setting of our experiments. The source code, data, and scripts will all be publicly available. Our findings are in alignment with observed empirical outcomes.
Appendix A Evaluation Metrics
The memorization accuracy (MA) is defined as follows:
$MA=\frac{1}{T}\sum_{t=1}^{T}\mathcal{A}_{t},$ (1)
where $T$ is the number of incremental steps. $\mathcal{A}_{t}$ represents average accuracy on the training instances from all learned concepts. Specifically, $\mathcal{A}_{t}$ is computed as follows:
$\mathcal{A}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (2)
where $a_{t,i}$ represents the accuracy evaluated on the task $i$ after training the model incrementally from concepts belonging to group $1,\cdots,t$. The accuracy is calculated as the exact match between the model output and the target output.
The memorization forgetting (MF) is computed as the average accuracy on all training instances of all learned concepts:
$MF=\frac{1}{T-1}\Sigma_{i=1}^{T-1}[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}].$ (3)
Concretely, $max_{j<T}(\{a_{j,i}\}_{j})$ represents the highest accuracy of task $i$ since it has been learned. $a_{T,i}$ represents the accuracy of task $i$ at step $T$. $[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}]$ computes the decrease of the accuracy of task $i$ at incremental step $T$.
The generalization accuracy (GA) and generalization forgetting (GF) are computed in a similar way except that the model is evaluated on the test set instead of the training set.
Table 6: The statistics of the PLMs used in this paper. $\dagger$: Non-embedding parameters according to Biderman et al. (2023).
Model Class Pretrained Weights Parameters Layers Hidden Dim Link
GPT-NeoX Pythia-70m 19M 6 512 Link
Pythia-160m 85M 12 768 Link
Pythia-410m 302M 24 1024 Link
Pythia-1b 805M 16 2048 Link
Pythia-1.4b 1.21B 24 2048 Link
Pythia-2.8b 2.52B 32 2560 Link
LLaMa llama-7b-hf 7B 32 4096 Link
vicuna-7b-v1.1 7B 32 4096 Link
llama-2-13b-hf 13B 40 5120 Link
Appendix B Experimental Settings
Backbones. We use the Pythia suite Biderman et al. (2023) and the popular open-source models including LLaMa and Vicuna for experiments. Pythia is based on GPT-NeoX Black et al. (2022), which contains 8 model sizes and 154 pre-training checkpoints, enabling research in interpretability and learning dynamics. The statistics of the 9 PLMs used in this paper are summarised in Table 6. We download the pre-trained weights from Huggingface (Wolf et al., 2019).
Datasets. We split the 1023 concepts in Concept-1K into 10 groups for IIL. Specifically, they sorted the concepts alphabetically and shuffled the order with random seed 1. The first group contains 105 concepts, while the others contain 102 concepts.
Baseline Methods. The introduction of the baseline methods is as follows:
  • SEQ: SEQ refers to sequential fine-tuning. It can be considered as the lower bound of incremental learning.
  • REPLAY: Experience replay stores representative old samples and jointly optimizes old and new samples when learning new tasks. It is a practical and popular technique in incremental learning.
  • EWC Kirkpatrick et al. (2017): EWC is a regularization-based method. The weight of each parameter is determined by the diagonal of the Fisher information matrix. The weight of the regularization loss is searched in {5×$\times 10^{3}$, 1×$\times 10^{4}$, 5×$\times 10^{4}$, 1×$\times 10^{5}$, 1×$\times 10^{6}$, 1×$\times 10^{7}$}.
  • LAMOL Sun et al. (2020): LAMOL trains PLMs with question-answering and generative objectives. It generates pseudo-samples before learning each new task for data replay. The weight of the generation loss is $\lambda=0.25$. The proportion of pseudo samples is $\gamma=0.20$. There are two variants of LAMOL, including LAMOL_t and LAMOL_g. The difference is whether a task-specific token is used for generation.
  • L2KD Chuang et al. (2020): L2KD adds a knowledge distillation target based on LAMOL. The teacher model is trained from scratch. We implemented the word-level variant since it performs best on text classification tasks.
  • LAMOL_KD Zheng et al. (2023): LAMOL_KD utilizes knowledge distillation based on LAMOL_t. Unlike L2KD, the teacher model in LAMOL_KD is trained on all previous tasks. The new data are used to learn the LAMOL objectives, and pseudo data are used for word-level knowledge distillation as a regularization term.
  • PCLL Zhao et al. (2022): PCLL combines the idea of variational autoencoders and word-level knowledge distillation with the objective of LAMOL.
  • LFPT5 Qin and Joty (2022): LFPT5 only learns soft prompts for each new task. The training objective is the same as LAMOL. The number of soft prompt tokens is 10.
  • LoRA Hu et al. (2021): LoRA only trains a small proportion of parameters of PLMs. We set the rank $r=8$ and the scaling parameter $\alpha=16$. We use the implementation of LoRA from the PEFT library Mangrulkar et al. (2022).
We note that some IL methods are not compared since they are not applicable in the IIL scenario. For example, Progressive Prompt Razdaibiedina et al. (2023) requires task ID during both training and inference stages. VAG Shao et al. (2023) requires storing the vocabulary of class labels and do not apply to the generation tasks without class labels. Besides, prompt-based IL methods such as L2P Wang et al. (2022) also do not apply to generation tasks.
Implementation Details. The maximum input and output lengths are set as 32 and 10, respectively. The batch size is 32. The learning rate of PLMs is $1\times 10^{-5}$. We use AdamW Loshchilov and Hutter (2018) optimizer. We use A800 GPUs for our experiments. Each experiment is repeated three times, and the average and standard deviations are reported. We search for the best hyper-parameters for each baseline method.
Figure 3: The accuracy of different concepts in the first task after sequential finetuning Pythia-2.8B on the 10 tasks of Concept-1K.
Appendix C Additional Experimental Results and Analysis
C.1 The Role of Model Scale, Pretraining, and Buffer Size
We provide the detailed result of Figure 2 in Table 7, 8, 9, 10, 11, 12, 13, and 14.
C.2 The performance of SOTA methods
The backbone model is Pythia-410M. We provide the step-wise average accuracy on training and test set in Figure 4 (a) and (b). The results demonstrate a large gap between SOTA IL methods and data replay. We also provided a comparison between SOTA methods in terms of memorization accuracy, generalization accuracy, memorization forgetting, and generalization forgetting in Figure 4 (c), (d), (e), and (f). In addition, from Figure 4 (g), we find that the training loss of the prompt-tuning-based method LFPT5 can not decrease to a small value. It indicates that only finetuning prompts are not practical for learning new knowledge, which is also consistent with the result of LoRA.
C.3 In-Context Learning
Apart from the analysis in the main paper, Table 5 shows that the smallest PLM (Pythia-70M) achieve high accuracy under the “1 shot” and “Same Know.” settings. The reason is that small PLMs simply copy the output in the demonstration instance as the final output. Under the “5 shot” and “Same Know.” settings, the accuracy of Pythia-70M is only 15.02%.
We use the model “gpt-3.5-turbo” and “gpt-4-turbo” for GPT3.5 and GPT4, respectively. Due to the high cost of evaluating on the whole concept-1K, we randomly sample 500 instances for GPT3.5 and GPT4. The output and the target of GPT3.5 and GPT4 on the 500 instances are provided in the supplementary material. Table 5 shows that GPT4 achieves 86.20% under the “5 shot” and “Same Know.” settings. It indicates that the training and the test instance of Concept-1K share the same knowledge.
C.4 Concept-Level Accuracy on Concept-1K
We analyze the memorization accuracy of the concepts in the first task after IIL. The backbone model is Pythia-2.8B, and it is sequential finetuning on Concept-1K. From Figure 3, we have the following observation:
  • The concepts such as “Cultural Competence”, “Augmented Reality Displays”, “Geopolitical Risks” that achieve high accuracy are broad, interdisciplinary topics that are likely rich in diverse data sources and content. These subjects span various fields and are likely to have clear definitions, significant public interest, and extensive coverage in the pretraining datasets, facilitating high memorization accuracy.
  • The lowest accuracy is observed in highly technical, cutting-edge, or niche fields such as “Spatial Audio”, “Biomaterials”, “Sustainable Tech”, and “Space Tourism”. These topics are likely challenging due to rapidly evolving information, specialized knowledge requirements, and potentially limited representation in available datasets.
In summary, the tendency from high to low accuracy reflects the transition from broad, interdisciplinary topics with abundant and diverse data to more specialized, technical, or emerging fields with potentially limited data exposure. This suggests that the PLMs’ memory accuracy is influenced by the richness and diversity of information available in its training data and the clarity and stability of the concepts themselves. Additionally, despite their significance, emerging technologies and areas of innovation might be harder for PLMs to learn accurately due to the rapid evolution of knowledge and possibly less structured data.
Table 7: The memorization accuracy when different model scales and buffer size are selected. The pretraining step is 143000 (final version). Figure 2 (a) summarises this figure’s content.
70M 160M 410M 1B 1.4B 2.8B
Buffer Size=0 6.34±0.86 33.53±1.39 58.28±0.61 65.97±0.31 64.95±0.55 68.03±0.46
Buffer Size=2000 22.87±1.28 56.12±0.81 77.31±0.23 81.57±0.32 80.81±0.62 82.26±0.66
Buffer Size=5000 42.13±1.88 73.32±1.00 90.05±0.64 91.93±0.49 91.49±0.32 92.03±0.35
Buffer Size=10000 58.50±0.50 84.91±1.64 97.81±0.12 97.21±0.73 97.35±0.84 96.74±0.43
Buffer Size=All 62.37±1.82 90.04±0.64 99.01±0.14 99.10±0.57 98.86±0.55 98.72±0.34
Table 8: The generalization accuracy when different model scales and buffer size are selected. The pretraining step is 143000 (final version). Figure 2 (e) summarises this figure’s content.
70M 160M 410M 1B 1.4B 2.8B
Buffer Size=0 2.52±0.12 6.57±0.28 17.69±0.35 22.65±0.39 25.57±0.36 29.56±0.35
Buffer Size=2000 4.04±0.07 8.49±0.23 22.49±0.32 27.29±0.55 30.80±0.28 34.21±0.40
Buffer Size=5000 5.14±0.09 10.08±0.40 25.19±0.46 29.34±0.38 33.50±0.46 37.04±0.36
Buffer Size=10000 5.88±0.10 10.36±0.52 26.20±0.35 30.81±0.24 34.85±0.57 37.72±0.52
Buffer Size=All 5.80±0.08 10.80±0.46 25.70±0.28 30.99±0.42 35.03±0.27 37.97±0.37
Table 9: The memorization accuracy when different pretraining steps and model scales are selected. The buffer size is 0. Figure 2 (b) summarizes the content of this figure.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 2.92±0.06 4.89±0.16 4.58±0.10 24.69±0.04 36.08±0.63 6.34±0.86
Pythia-160M 26.33±0.31 29.34±0.25 29.11±0.11 48.78±0.09 60.12±0.34 33.53±1.39
Pythia-410M 29.58±1.10 31.06±0.24 30.78±0.22 49.84±0.06 64.31±0.38 58.28±0.61
Pythia-1B 33.18±0.59 33.71±0.17 33.91±0.49 53.14±0.59 66.66±0.59 65.97±0.31
Pythia-1.4B 35.61±0.59 36.31±0.51 36.15±0.32 55.61±0.59 70.66±0.51 64.95±0.55
Pythia-2.8B 37.76±0.29 37.73±0.38 37.65±0.39 57.09±0.34 72.92±0.59 68.03±0.46
Table 10: The generalization accuracy when different pretraining steps and model scales are selected. The buffer size is 0. Figure 2 (f) summarises this figure’s content.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 0.83±0.02 1.10±0.01 1.08±0.04 2.94±0.04 3.95±0.03 2.52±0.12
Pythia-160M 1.64±0.04 1.81±0.02 1.82±0.06 3.79±0.05 6.96±0.04 6.57±0.28
Pythia-410M 2.34±0.07 2.35±0.04 2.21±0.04 4.18±0.10 12.91±0.14 17.69±0.35
Pythia-1B 2.78±0.14 2.85±0.24 2.80±0.24 6.03±0.24 17.70±0.24 22.65±0.39
Pythia-1.4B 2.71±0.37 2.32±0.24 2.40±0.53 5.68±0.24 20.64±0.24 25.57±0.36
Pythia-2.8B 2.48±0.23 1.79±0.29 2.09±0.32 5.35±0.49 24.79±0.24 29.56±0.35
Table 11: The memorization accuracy when different pretraining steps and model scales are selected. The buffer size is 2000. Figure 2 (c) summarises this figure’s content.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 22.15±0.13 24.91±0.20 23.54±0.15 55.06±0.16 64.89±0.23 22.87±1.28
Pythia-160M 59.42±0.13 59.62±0.77 59.30±0.18 78.84±0.44 82.30±0.18 56.12±0.81
Pythia-410M 60.43±0.59 61.09±0.66 61.24±0.27 79.77±0.15 83.24±0.12 77.31±0.23
Pythia-1B 63.11±0.15 62.05±0.40 63.51±0.31 80.11±0.50 83.11±0.38 81.57±0.32
Pythia-1.4B 63.77±0.15 64.14±0.23 65.40±0.32 82.38±0.34 84.19±0.22 80.81±0.62
Pythia-2.8B 65.10±0.45 65.84±0.57 66.04±0.53 82.63±0.18 84.80±0.44 82.26±0.66
Table 12: The generalization accuracy when different pretraining steps and model scales are selected. The buffer size is 2000. Figure 2 (g) summarises this figure’s content.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 2.07±0.08 2.46±0.06 2.51±0.13 3.84±0.06 4.57±0.07 4.04±0.07
Pythia-160M 3.17±0.08 3.55±0.14 3.52±0.06 5.26±0.14 8.05±0.08 8.49±0.23
Pythia-410M 3.96±0.06 4.97±0.09 5.09±0.04 6.25±0.13 15.64±0.06 22.49±0.32
Pythia-1B 5.51±0.17 5.81±0.43 6.37±0.56 8.85±0.55 20.01±0.51 27.29±0.55
Pythia-1.4B 6.76±0.21 6.98±0.16 7.55±0.55 10.21±0.42 23.67±0.18 30.80±0.28
Pythia-2.8B 8.07±0.28 8.57±0.54 9.32±0.40 12.61±0.25 27.50±0.23 34.21±0.40
Table 13: The memorization accuracy when different pretraining steps and model scales are selected. The buffer size is unlimited (all old samples are stored). Figure 2 (d) summarizes the content of this figure.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 65.77±0.56 67.75±0.19 66.36±0.09 87.61±0.14 92.59±0.11 62.37±1.82
Pythia-160M 86.54±0.15 88.75±0.28 88.64±0.07 98.95±0.03 99.62±0.05 90.04±0.64
Pythia-410M 87.46±0.19 89.07±0.27 90.03±0.16 99.30±0.22 99.59±0.02 99.01±0.14
Pythia-1B 88.77±0.02 89.75±0.19 90.36±0.09 99.75±0.19 99.36±0.09 99.10±0.57
Pythia-1.4B 89.65±0.02 90.34±0.03 90.76±0.54 99.85±0.21 99.76±0.02 98.86±0.55
Pythia-2.8B 90.31±0.36 90.89±0.53 91.43±0.44 99.95±0.56 99.86±0.02 98.72±0.34
Table 14: The generalization accuracy when different pretraining steps and model scales are selected. The buffer size is unlimited (all old samples are stored). Figure 2 (h) summarises this figure’s content.
Step=0 Step=16 Step=128 Step=1000 Step=10000 Step=143000
Pythia-70M 3.18±0.12 3.10±0.02 3.08±0.04 4.37±0.06 5.40±0.03 5.80±0.08
Pythia-160M 3.79±0.25 3.95±0.10 3.93±0.04 6.09±0.06 9.08±0.06 10.80±0.46
Pythia-410M 4.48±0.03 5.12±0.18 5.24±0.16 7.29±0.16 18.14±0.11 25.70±0.28
Pythia-1B 6.12±0.56 6.12±0.41 6.75±0.20 9.74±0.26 25.07±0.38 30.99±0.42
Pythia-1.4B 6.96±0.36 6.87±0.55 7.45±0.50 11.07±0.40 27.93±0.56 35.03±0.27
Pythia-2.8B 8.47±0.20 8.81±0.31 9.82±0.45 14.73±0.15 30.65±0.21 37.97±0.37
(a) Average Accuracy on Training Set
(b) Average Accuracy on Test Set
(c) Memorization Accuracy
(d) Generalization Accuracy
(e) Memorization Forgetting
(f) Generalization Forgetting
(g) Training Loss
Figure 4: The detailed result of SOTA methods on Concept-1K.
Table 15: Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance.
Domain Concept Triplet Training and Test Input Target Output
Science and Technology Brain-Computer Interface (Brain-Computer Interface, UsedFor, ControllingComputersWithThought) What is the main use of a Brain-Computer Interface? controlling computers with thought
How can individuals utilize a Brain-Computer Interface?
(Brain-Computer Interface, Uses, EEG) What does a Brain-Computer Interface use for its functioning? EEG
What technology is utilized by a Brain-Computer Interface?
(Brain-Computer Interface, DesignedFor, PersonsWithDisabilities) Who is the Brain-Computer Interface designed for? persons with disabilities
What group benefits from the Brain-Computer Interface’s design?
(Brain-Computer Interface, PartOf, Neurotechnology) What field is Brain-Computer Interface a part of? neurotechnology
To which broader technology does Brain-Computer Interface belong?
(Brain-Computer Interface, Requires, BrainSignals) What does a Brain-Computer Interface require to function? brain signals
What is necessary for the operation of a Brain-Computer Interface?
IPTV (IPTV, IsA, TelevisionService) What type of service is IPTV classified as? television service
How is IPTV categorized in terms of service?
(IPTV, CompetesWith, CableTelevision) Who does IPTV compete with? cable television
What is IPTV’s competition in the television market?
(IPTV, Supports, HighDefinition) What type of video quality is supported by IPTV? high definition
What does IPTV support?
(IPTV, Allows, TimeShifting) What does IPTV allow? time shifting
What feature does IPTV provide concerning watching schedules?
(IPTV, Needs, SetTopBox) What does IPTV need to operate? set top box
What equipment is necessary for IPTV functionality?
Education MOOCs (MOOCs, PartOf, OnlineLearning) What is MOOCs a part of? online learning
Under what broader learning category do MOOCs fall?
(MOOCs, AffiliatedWith, Universities) With whom are MOOCs affiliated? universities
What types of institutions are MOOCs commonly associated with?
(MOOCs, Offers, FreeCourses) What do MOOCs offer to learners? free courses
What can learners benefit from MOOCs at no cost?
(MOOCs, AimedAt, SelfLearners) Who are MOOCs primarily aimed at? self learners
What target audience do MOOCs cater to?
(MOOCs, Employs, AutomatedGrading) What grading method do MOOCs employ? automated grading
How are assignments typically graded in MOOCs?
STEM Education (STEM Education, FocusOn, ScienceTechnologyEngineeringMathematics) What does STEM Education focus on? science technology engineering mathematics
What are the main areas of study in STEM Education?
(STEM Education, Encourages, CriticalThinking) What does STEM Education encourage in students? critical thinking
What skill is STEM Education known to foster among learners?
(STEM Education, BenefitsFrom, InterdisciplinaryApproach) What does STEM Education benefit from? interdisciplinary approach
Which approach enhances the effectiveness of STEM Education?
(STEM Education, Enhances, TechnologicalLiteracy) What does STEM Education enhance among students? technological literacy
What aspect of literacy is improved through STEM Education?
(STEM Education, SupportedBy, GovernmentPolicies) Who supports STEM Education through policies? government policies
What supports the implementation of STEM Education?
Economy (Ethereum 2.0, Offers, Scalability) What does Ethereum 2.0 offer to its users? scalability
What is a key feature of Ethereum 2.0?
(Ethereum 2.0, Implements, ProofOfStake) What consensus mechanism does Ethereum 2.0 implement? proof of stake
Which algorithm is used by Ethereum 2.0 for transaction validation?
(Ethereum 2.0, AimsTo, ReduceEnergyConsumption) What is Ethereum 2.0’s goal regarding energy use? reduce energy consumption
How does Ethereum 2.0 plan to impact the environment?
(Ethereum 2.0, Uses, BeaconChain) What does Ethereum 2.0 use to manage its consensus process? beacon chain
What is a key component of Ethereum 2.0’s infrastructure?
(Ethereum 2.0, PlansFor, eWASM) What does Ethereum 2.0 plan for? eWASM
What technology is Ethereum 2.0 planning to incorporate?
Bitcoin (Bitcoin, CreatedBy, SatoshiNakamoto) Who created Bitcoin? satoshi nakamoto
By whom was Bitcoin developed?
(Bitcoin, Uses, BlockchainTechnology) What technology does Bitcoin utilize? blockchain technology
What is the basis of Bitcoin’s operation?
(Bitcoin, CapableOf, PeerToPeerTransactions) What type of transactions is Bitcoin capable of? peer-to-peer transactions
What kind of financial transactions can Bitcoin perform?
(Bitcoin, HasProperty, Decentralized) What property does Bitcoin have? decentralized
How is Bitcoin characterized in terms of its structure?
(Bitcoin, Causes, Speculation) What does Bitcoin cause in the market? speculation
What outcome is often associated with Bitcoin’s presence in the market?
Culture Vlogging (Vlogging, HasPrerequisite, ContentPlanning) What is a prerequisite for Vlogging? content planning
Before starting to Vlog, what must one prepare?
(Vlogging, CausesDesire, ToBeFamous) What desire does Vlogging cause? to be famous
Why do many people start Vlogging?
(Vlogging, Entails, VideoEditing) What does Vlogging entail? video editing
Besides shooting content, what is an essential skill for Vlogging?
(Vlogging, Causes, ViewerEngagement) What does vlogging increase? viewer engagement
What is a result of vlogging?
(Vlogging, HasLastSubevent, Uploading) What is the final stage in the process of Vlogging? uploading
What activity typically concludes the vlogging process?
Fan Fiction (Fan Fiction, IsA, LiteraryGenre) What is Fan Fiction classified as? literary genre
Under which category does Fan Fiction fall?
(Fan Fiction, InspiredBy, ExistingWorks) What inspires Fan Fiction? existing works
What is the source material for Fan Fiction?
(Fan Fiction, IsFor, Entertainment) What is the purpose of Fan Fiction? entertainment
Why do people engage with Fan Fiction?
(Fan Fiction, MayViolate, CopyrightLaw) What law may fan fiction violate? copyright law
What legal issue does fan fiction potentially encounter?
(Fan Fiction, Addresses, UnexploredPlotlines) What does Fan Fiction specifically address? unexplored plotlines
What kind of plotlines does Fan Fiction focus on?
Health and Medical Nanomedicine (Nanomedicine, UsedFor, Targeted Drug Delivery) What is Nanomedicine used for? targeted drug delivery
What application of Nanomedicine allows for precise medication delivery?
(Nanomedicine, MotivatedByGoal, Improve Treatment Efficacy) What goal motivates the field of nanomedicine? improve treatment efficacy
Why is nanomedicine being developed and researched?
(Nanomedicine, CapableOf, Crossing Blood-Brain Barrier) What is nanomedicine capable of in terms of biological barriers? crossing blood-brain barrier
How does nanomedicine interact with the blood-brain barrier?
(Nanomedicine, CausesDesire, Minimally Invasive Treatment) What does nanomedicine cause a desire for? minimally invasive treatment
What kind of treatment does nanomedicine make more desirable?
(Nanomedicine, Requires, Nanoparticles) What does Nanomedicine require to function? nanoparticles
What are the key components essential for Nanomedicine?
Nutritional Genomics (Nutritional Genomics, Studies, Gene-Nutrient Interaction) What does Nutritional Genomics primarily study? gene-nutrient interaction
What is the focus of research in Nutritional Genomics?
(Nutritional Genomics, AimsTo, Personalize Nutrition) What is the goal of Nutritional Genomics? personalize nutrition
What does Nutritional Genomics strive to achieve in its application?
(Nutritional Genomics, HelpsIn, Disease Prevention) How does Nutritional Genomics contribute to health? disease prevention
In what way does Nutritional Genomics play a role in healthcare?
(Nutritional Genomics, ContributesTo, Healthier Diet Planning) What does Nutritional Genomics contribute to? healthier diet planning
What aspect of health is Nutritional Genomics associated with improving?
(Nutritional Genomics, Affects, Nutrient Metabolism) What does Nutritional Genomics affect? nutrient metabolism
What aspect of health does Nutritional Genomics have an impact on?
Table 16: (Continual) Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance.
Domain Concept Triplet Training and Test Input Target Output
Environment Groundwater Recharge (Groundwater Recharge, IsA, HydrologicalProcess) What is Groundwater Recharge classified as? hydrological process
What kind of process is Groundwater Recharge?
(Groundwater Recharge, UsedFor, AquiferSustainability) What is Groundwater Recharge used for? aquifer sustainability
What purpose does Groundwater Recharge serve in relation to aquifers?
(Groundwater Recharge, Requires, PermeableSurfaces) What does Groundwater Recharge require? permeable surfaces
What are essential for the process of Groundwater Recharge?
(Groundwater Recharge, ResultsIn, WaterTableRise) What is a result of Groundwater Recharge? water table rise
What does Groundwater Recharge lead to regarding water tables?
(Groundwater Recharge, MotivatedByGoal, DroughtMitigation) What goal motivates groundwater recharge? drought mitigation
Why is groundwater recharge important?
Sea Level Rise (Sea Level Rise, CausedBy, GlobalWarming) What causes Sea Level Rise? global warming
What is the primary factor leading to Sea Level Rise?
(Sea Level Rise, AnalyzedBy, Climatologists) Who analyzes Sea Level Rise? climatologists
What group of professionals study Sea Level Rise?
(Sea Level Rise, ResultsIn, HabitatLoss) What does Sea Level Rise result in? habitat loss
What is a significant impact of Sea Level Rise on natural habitats?
(Sea Level Rise, MeasuredBy, TideGauges) How is sea level rise measured? tide gauges
What instrument is used to measure sea level rise?
(Sea Level Rise, AddressedBy, EmissionReductions) How is Sea Level Rise addressed? emission reductions
What strategy addresses Sea Level Rise?
Figure 5: The histogram of the relations in Concept-1K. The top-50 relations are shown.
Figure 6: The word cloud diagram of Concept-1K.
Table 17: The concept list of Concept-1K. The concepts are sorted in alphabetically order.
360-Degree Videos Bioprinting Organs Critical Thinking Documentary Collections
3D Audio Biotechnology Cross-border e-Commerce Documentary Films
3D Bioprinting Bitcoin Cross-border E-commerce Drone Technology
3D Modeling Bitcoin Halving Cross-Border Healthcare Drought Resilience
3D Printed Drugs Blended Learning Cross-Border Payments Drug Delivery Systems
3D Printing Blockchain Health Records Cross-Cultural Exchange Dynamic Game Content
3D Rendering Blockchain Supply Chain Cross-Platform Gaming E-Book Popularity
3D Rendering Engines Blockchain Technology Crowdfunding E-Books in Education
3D Scanning Brain-Computer Interface Crowdfunding Arts Eco-Cities
5G Networks Branding Crowdfunding Innovations Eco-Friendly
5G Technology Brexit Impact Crowdfunding Platforms Eco-Friendly Medical Products
Academic Journals Online Broadcast Graphics Crowdsourced Projects E-commerce
Accessibility in Design Broadcast Technology Cryptocurrency e-Commerce Innovation
Acoustic Modeling Bug Bounty Programs Cryptocurrency Adoption E-Commerce Technology
Activity Trackers Business Model Innovation Cryptocurrency Exchanges Economic Sanctions
Adaptive Biotechnologies Business Process Outsourcing Cryptocurrency Investing Ecosystem Services
Adaptive Game AI Cable TV Technology Cryptocurrency Mining Ecotourism
Adaptive Learning California Consumer Privacy Act Cryptography Eco-Tourism
Addiction Recovery Apps Cancer Immunotherapy Cultural Competence Edge Computing
Additive Manufacturing Car Sharing Cultured Meat Edtech Evolution
Adult Education Carbon Capture Currency Fluctuations EdTech Innovation
Affiliate Marketing Carbon Credits Currency Hedging Educational Apps
Agile Methodology Carbon Footprint Customer Experience Educational Data Mining
Agile UX Carbon Neutral Customer Relationship Management Educational Games
Agri-Robots Carbon Tax Cyber Activism Educational Podcasts
AgriTech CBDC Cyber Forensics Educational Vlogging
Agroforestry Central Bank Policies Cybernetic Implants Edutainment
Agtech Innovation CGI Cybersecurity E-Governance
AI Diagnostics CGI Animation Cybersecurity Investments E-Health
AI Drug Discovery Challenger Banks DAOs E-Learning
AI Epidemiology Chatbot Design Dark Matter Research Electric Vehicles
AI Ethics Chatbot Technology Data Privacy Regulations Electronic Health Records
AI Governance Chatbots Data Science Electronic Skin Patches
AI in Education Cinema Podcasts Data Security Email Marketing
AI in Radiology Cinematography Decentralized Finance E-Mentoring
AI Pathology Circular Economy Deep Learning Emerging Markets
Air Pollution Sensors Citizen Journalism DeFi Platforms Emerging Markets Growth
Algorithmic Trading Civic Tech Deflation Risks Encryption Algorithms
Alternative Credentialing Classroom Technologies Deforestation Endangered Species
Alternative Proteins Clean Technology Dementia Care Technology Energy Efficiency
Animation Advances Climate Adaptation Design Sprints Environmental Advocacy
Animation in UI Climate Change Design Thinking Environmental Economics
Animation Software Climate Finance DevOps Practices Environmental Education
Animation Technology Climate Legislation Diet Tracking Apps Environmental Health Surveillance
Antibiotic Stewardship Climate Mitigation Digital Advertising Environmental Monitoring
Anti-Money Laundering Solutions Climate Tech Digital Art Environmental Policy
Antitrust Regulations Clinical Decision Support Digital Asset Management Environmental Protection
Antiviral Therapies Cloud Computing Digital Badges Environmental Sustainability
API Economy Cloud Gaming Digital Broadcasting E-Paper Technology
Apprenticeships Cloud Storage Digital Cinema E-Portfolios
Aquaponics Cloud Video Editing Digital Cinematography ESG Investing
AR Contact Lenses Cloud-Based PACS Digital Collaboration ESG Reporting
Artificial Intelligence Coastal Erosion Digital Comics ESL/EFL Teaching
Astrophysics Coding Education Digital Content Creation eSports
Attendance Tracking Cognitive Behavioral Apps Digital Creators eSports Growth
Audience Measurement Cognitive Walkthrough Digital Currencies eSports Platforms
Audio Books Collaborative Editing Digital Curriculum eSports Tournaments
Audio Compression Collaborative Learning Digital Education Ethereum
Audio Editing Software Collaborative Platforms Digital Festivals Ethereum 2.0
Audio Recognition Color Grading Digital Health Ethical Clothing
Audio Synthesis Color Theory Digital Illustration Ethical Hacking
Augmented Reality Commodity Markets Digital Journalism EV Battery Technology
Augmented Reality Content Community Learning Digital Libraries Exercise Apps
Augmented Reality Displays Competency-Based Education Digital Literacy Exit Strategies
Augmented Reality Experiences Composting Digital Marketing Exoplanet Discovery
Augmented Reality Games Computational Thinking Digital Mental Health Experiential Learning
Augmented Reality Gaming Computer Hardware Digital Nomadism Export Controls
Augmented Reality Learning Computer Vision Digital Patient Engagement Export Credit Insurance
Augmented Reality Retail Computer-aided Design Digital Pharmacy Export Import
Augmented Reality Surgery Connected Devices Digital Portfolios FaaS
Automated Dispensing Conservation Efforts Digital Psychiatry Facial Recognition
Autonomous Vehicles Contactless Payments Digital Rights Facial Recognition Tech
BaaS Containerization Digital Rights Management Fact-Checking
Backend Development Content Delivery Networks Digital Runways Fan Fiction
Battery Technologies Content Distribution Networks Digital Sculpture Farm Management Software
Beekeeping Content Management Systems Digital Signal Processing Fashion Apps
Behavioral Health Integration Content Marketing Digital Storytelling Fashion Blogging
Big Data Continuous Assessment Digital Textbooks Feedback Culture
Big Data Analytics Continuous Glucose Monitors Digital Therapeutics Fermented Foods
Bike Sharing Continuous Patient Monitoring Digital Transformation Field Trips
Bilingual Education Conversational AI Digital Twins Film Festivals
Biodiversity Conversational Design Digital Twins Healthcare Film Production Tech
Bioelectronic Medicine Corporate Restructuring Digital Wallets Film Reviews
Bioenergy Corporate Sustainability Digital Yuan Film Scoring
Biohacking Counseling Services Direct Listings Film Workshops
Biomaterials Course Management Systems Disease Prediction AI Financial Crime Compliance
Biomedical Engineering COVID-19 Pandemic Distance Learning Financial Inclusion
Biometric Authentication Credit Scoring Distributed Computing Financial Stability
Biometric Monitoring CRISPR Cas-9 Diversity Training Financial Technology
Biometric Systems CRISPR Technology Docker Technology Fintech Evolution
Table 18: (Continual) The concept list of Concept-1K. The concepts are sorted in alphabetically order.
Firewall Technologies Health Analytics Live Streaming Network Monitoring
Fiscal Stimulus Health Coaching Bots Live Video Technology Network Security
Fitness Technology Health Data Analytics Livestream Shopping Net-Zero Targets
Fitness Trackers Health Data Exchange Localization in Gaming Neural Networks
Fitness Wearables Health Data Privacy Logistics Technology Neurofeedback
Flexible Displays Health Equity Longevity Medicine Neuromodulation
Flipped Classroom Health Equity Solutions Low-Carbon Technology News Podcasts
Flood Management Health Gamification Machine Learning NFT Art
Food Safety Technologies Health Informatics Machine Learning Teaching NFT Marketplaces
Food Security Health IoT Malware Analysis NFTs
Food Technology Health Literacy Platforms Manufacturing Reshoring Non-Invasive Diagnostics
Food Traceability Health Social Networks Market Regulation Changes Nutraceuticals
Foreign Exchange Healthtech Advances Market Volatility Analysis Nutrigenomics
Forex Market Trends Herbicides Marketing Automation Nutritional Genomics
Fraud Detection Heuristic Evaluation Marketing Strategies Nutritional Tech
Freemium Gaming Models High-Frequency Trading Mars Missions Ocean Acidification
Frontend Development High-frequency Trading Mastery Learning Ocean Economy
Full-stack Development High-performance Computing Material Science Offshoring
Functional Foods Holistic Health Approaches Media Analytics Oil Price Dynamics
Game Accessibility Features Home Workout Solutions Media Encoding OLED Technology
Game AI Human-Centered Design Media Literacy Omni-channel Retailing
Game Development Engines Human-Computer Interaction Media Monitoring Online Assessments
Game Development Tools Humanoid Robots Media Storage Solutions Online Book Clubs
Game Engines Hybrid Vehicles Media Transcoding Online Certifications
Game Level Design Hydroelectric Power Medical Chatbots Online Communities
Game Monetization Models Hydroponics Medical Devices Online Conferences
Game Optimization IaaS Medical Drones Online Courses
Game Sound Design ICOs Medical Imaging AI Online Courseware
Game Streaming Identity Management Medical Tricorders Online Education
Game Voice Acting IIoT Medical Wearables Online Fairs
Game World Building Immersive Storytelling Mental Health Apps Online Gaming Infrastructure
Gamification in Education Immune System Mapping Mental Health Awareness Online Learning Platforms
Gamification Techniques Immunogenomics Mental Health Chatbots Online Museums
Gaming Consoles Immunotherapy Advances Mental Health Monitoring Online Newsrooms
Gaming Conventions Impact Investing Mental Health Technologies Online Petitions
Gaming Headsets Inclusive Education Mental Wellness Apps Online Retail
Gaming Influencers Inclusive Game Design Mergers and Acquisitions Online Styling
Gaming Monetization Independent Filmmakers Metamaterials Online Tutoring
Gaming Peripherals Industrial Automation Metaverse Online Workshops
Gaming Platforms Industrial Design Microbiome Analysis Open Access Publishing
Gaming SDKs Industrial Robots Micro-Credentials Open Banking
GDPR Compliance Industry 4.0 Microfinance Open Educational Resources
Gender Lens Investing Inflation Trends Microfinance Growth Open Source Software
Gene Editing Influencer Culture Microgravity Research Opioid Alternatives
Gene Therapy Influencer Marketing Micro-interactions Orbital Mechanics
Genetic Counseling Infographics Microloans Organic Farming
Genetic Risk Assessment Information Architecture Microprocessors Organ-on-a-Chip
Genetic Testing In-game Advertising Microservices OTT Services
Genomic Sequencing In-Game Advertising Mindfulness Apps Outcome-Based Education
Genomics In-Game Purchases Mixed Reality Outdoor Education
Geopolitical Risks Insurtech Mixed Reality Applications Outsourcing
Geopolitical Tensions Insurtech Trends Mixed Reality Classrooms P2P
Geothermal Energy Integrated Care Models Mixed Reality Gaming PaaS
Gestural Interfaces Intelligent Traffic Systems Mobile Applications Pandemic Preparedness
Gesture Control Interaction Design Mobile Banking Patient Empowerment
Gig Economy Interactive Exhibits Mobile Commerce Payment Gateways
Gigabit Society Interactive Game Design Mobile Game Development Payment Processing
Global Citizenship Interactive Storytelling Mobile Gaming Pay-per-click Advertising
Global Economic Recovery Interactive Whiteboards Mobile Gaming Hardware Pay-Per-View Services
Global Education Interest Rate Forecasts Mobile Health Peer Tutoring
Global Health Security Interest Rate Hikes Mobile Health Applications Peer-to-Peer Lending
Global Sourcing International Baccalaureate Mobile Health Clinics Penetration Testing
Global Supply Chains International Marketing Mobile Learning Performance Analytics
Global Trade Tensions International Payments Mobile Medical Units Permaculture
Global Warming International Students Mobile Payments Persona Development
Government Bonds International Trade Mobility as a Service Personal Assistants
Graphene Applications Internships Mockups Personal Health Analytics
Graphic Design Software Intrusion Detection MOOCs Personal Health Records
GraphQL Invasive Species Motion Capture Technology Personalized Learning
Green Bonds Inventory Management Motion Control Gaming Personalized Medicine
Green Building Materials IoT Motion Graphics Personalized Nutrition
Green Corridors IoT Security Motion Graphics Design Pesticides
Green Finance IPOs Movie Streaming Pharmacogenetics
Green Infrastructure IPTV mRNA Vaccines Pharmacogenomics
Green Living Know Your Customer Multi-camera Setup Photodynamic Therapy
Green Roofs Knowledge Process Outsourcing Multi-factor Authentication Photonic Crystals
Green Schools Kubernetes Multilingual Education Physics Engines
Green Spaces Landscape Ecology Multimedia Production Planetary Science
Green Technology Language Learning Apps Multiplayer Game Servers Plant-Based Diets
Greenhouse Gases Lean UX Music Recommendation Systems Plant-based Proteins
Grid Computing Learning Analytics Nanomedicine Plastic Alternatives
Groundwater Recharge Learning Management Systems Nanotech Drug Delivery Plastic Pollution
Group Projects Learning Platforms Native Species Player Behavior Analysis
Gut Microbiome LEED Certification Natural Capital Player Engagement Metrics
Habitat Destruction Letters of Credit Natural Language Processing Podcast Popularity
Hacking Lifelong Learning Nature Connectivity Podcasting
Handheld Gaming Devices Liquidity Mining Nearshoring Podcasting Tech
Haptic Feedback Devices Literary Podcasts Neobanks Podcasts in Education
Haptic Technology Live Broadcasting Neobanks Emergence Poetry Slams
Hashtag Campaigns Live Game Streaming Net Zero Pollinator Conservation
Table 19: (Continual) The concept list of Concept-1K. The concepts are sorted in alphabetically order.
Pollution Prevention Smart Manufacturing Toxic Chemicals Wearable Computers
Population Health Management Smart Materials Trade Agreements Wearable Fitness Trackers
Post-production Workflow Smart Meters Trade Finance Wearable Gaming Devices
Precious Metals Investments Smart Prosthetics Trade Wars Wearable Health
Precision Agriculture Smart Speakers Traffic Management Wearable Health Devices
Precision Farming Smart Sutures Typography Wearable Medical Technology
Precision Medicine Smart Transportation UI Design Patterns Wearable Sports Technology
Predictive Analytics Smart Waste Management Unicorn Startups Wearable Tech
Predictive Healthcare Analytics Smart Watches Upcycling Web Analytics
Printmaking Social Commerce Urban Analytics Web Development
Privacy Protection Social Gaming Urban Farming Web Novels
Private Equity Trends Social Impact Bonds Urban Farming Solutions Web3
Probiotics Social Learning Urban Forestry Webcasting
Procedural Generation Social Media Marketing Urban Greening Webinars
Product Design Social Media Trends Urban Mobility Wellness Apps
Professional Development Social Movements Urban Sustainability Wellness Economy
Protectionism Social News Urban Wildlife Wellness Programs
Prototype Design Socially Responsible Investing Usability Testing Wetland Restoration
Public Transportation Software Development User Experience Design Wildlife Conservation
Quantitative Trading Soil Contamination User Experience in Gaming Wildlife Tourism
Quantum Computing Solar Energy Finance User Flow Diagrams Wind Energy
Quantum Cryptography Solar Power User Interface Design Wind Energy Investments
Quantum Dots Sound Analysis User Journey Mapping Wireframing
Quantum Mechanics Sound Engineering User Research Wireless Networking
Racial Equity Investing Sound Mixing UX Metrics Work from Home
Radiomics Sovereign Debt Issues UX Writing XaaS
Rainwater Harvesting Space Exploration UX/UI Design XR
Ransomware Space Telescopes Vaccine Technology XR Gaming Experiences
Recycling Space Tourism Venture Capital Shifts Yield Farming
Regenerative Medicine Spatial Audio Vertical Farming YouTube Creators
Regtech Speech Synthesis Video Content Creation Zero Trust Security
Regtech Solutions Speech-to-Text Video Marketing Zero Waste
Reinforcement Learning Sports Data Analysis Video on Demand Zero Waste Lifestyle
Remote Health Consultations Sports Tech Video Streaming
Remote ICU Monitoring Stagflation Concerns Viral Challenges
Remote Learning Stakeholder Capitalism Virtual Assets
Remote Monitoring Tools Startup Ecosystems Virtual Assistants
Remote Patient Monitoring STEAM Education Virtual Book Tours
Remote Performances Stem Cell Research Virtual Care
Remote Therapy Sessions Stem Cell Therapy Virtual Classrooms
Remote Video Production STEM Education Virtual Clinical Trials
Remote Work Technologies Stormwater Management Virtual Concerts
Remote Workshops Storyboarding Virtual Conferences
Renewable Energy STOs Virtual Desktops
Responsible Travel Streaming Audio Virtual Events
Responsive Web Design Streaming Services Virtual Exhibitions
RESTful Services Stress Management Tools Virtual Fashion
River Restoration Student Engagement Virtual Fitting Rooms
Robo-advisors Student Engagement Tools Virtual Galleries
Robotic Surgery Student Wellbeing Virtual Goods
Robotics Study Abroad Programs Virtual Gyms
Robotics in Education Style Influencers Virtual Health Assistants
Rocket Science Subscription Gaming Services Virtual Health Fairs
SaaS Supercomputing Virtual Keyboards
Sales Automation Supply Chain Finance Virtual Protests
Satellite Broadcasting Supply Chain Innovation Virtual Prototyping
Satellite Internet Supply Chain Optimization Virtual Reality
Satellite Technology Sustainable Agriculture Virtual Reality Arcades
Scale-up Companies Sustainable Cities Virtual Reality Education
Screencasting Sustainable Development Virtual Reality Experiences
Screenwriting Sustainable Development Goals Virtual Reality Filmmaking
Scrum Framework Sustainable Energy Virtual Reality Gaming
Sea Level Rise Sustainable Fashion Virtual Reality Sports
Self-Publishing Sustainable Fishing Virtual Reality Theater
Semantic Analysis Sustainable Healthcare Virtual Reality Therapy
Semiconductor Tech Sustainable Investing Virtual Reality Training
SEO Sustainable Tech Virtual Set Design
Serious Games Sustainable Transport Virtual Shopping
Serverless Architecture Synthetic Biology Virtual Teaching Assistants
Service Design Tariff Implications Virtual Tours
Service Robots Tariff Negotiations Visual Effects
Sharing Economy Teacher Training Visual Effects Software
Shipping Solutions Telehealth Expansion Visual Hierarchy
Short Films Telehealth Licensing Visual Storytelling
Simulation Games Telehealth Solutions Vlogging
Skill Sharing Telemedicine Vocational Training
Sleep Tech Devices Teleophthalmology Voice Recognition
Sleep Technology Telepsychiatry Voice User Interface
Slow Fashion Tele-rehabilitation Vulnerability Assessment
Smart Beds Telestroke Services Walkability
Smart Cities Television Production Warehouse Automation
Smart Cities Technologies Text-to-Speech Waste Management
Smart Contracts Theoretical Physics Waste Reduction
Smart Diagnostic Wearables Therapeutic Games Water Conservation
Smart Glasses Therapeutic Robots Water Quality Monitoring
Smart Grids Threat Intelligence Water Scarcity Solutions
Smart Health Watches TikTok Dance Watershed Management
Smart Home Tissue Engineering Wealth Gap
Smart Inhalers Tokenization Wealthtech
Smart Lighting Tokenization of Assets Wealthtech Advancements

### v3

@@
-Concept-1K: A Novel Benchmark for Instance Incremental Learning
+Can LLMs Learn New Concepts Incrementally without Forgetting?
 Abstract
-Incremental learning (IL) is essential to realize the human-level intelligence in the neural network.
-However, existing IL scenarios and datasets are unqualified for assessing forgetting in PLMs, giving an illusion that PLMs do not suffer from catastrophic forgetting.
-To this end, we propose a challenging IL scenario called instance-incremental learning (IIL) and a novel dataset called Concept-1K, which supports an order of magnitude larger IL steps.
-Based on the experiments on Concept-1K, we reveal that billion-parameter PLMs still suffer from catastrophic forgetting, and the forgetting is affected by both model scale, pretraining, and buffer size.
-Furthermore, existing IL methods and a popular finetuning technique, LoRA, fail to achieve satisfactory performance.
-Our study provides a novel scenario for future studies to explore the catastrophic forgetting of PLMs and encourage more powerful techniques to be designed for alleviating the forgetting in PLMs.
+Large Language Models (LLMs) have achieved remarkable success across various tasks, yet their ability to learn incrementally without forgetting remains underexplored.
+Incremental learning (IL) is crucial as it enables models to acquire new knowledge while retaining previously learned information, akin to human learning.
+Existing benchmarks for IL are insufficient due to data leakage issues and the overqualification of LLMs.
+To address these challenges, we introduce Concept-1K, a novel dataset comprising 1,023 recently emerged concepts across diverse domains.
+The concepts in Concept-1K are discrete, interpretable units of knowledge that allow for fine-grained analysis of learning and forgetting processes.
+Using Concept-1K as a testbed, we aim to answer the question: “Can LLMs learn new concepts incrementally without forgetting like humans?” Our investigation reveals that LLMs still suffer from catastrophic forgetting and that LoRA, despite fine-tuning fewer parameters, may lead to more forgetting on training data.
+Additionally, we explore the roles of in-context learning, model scale, buffer size, and pretraining in IL performance.
+These findings highlight the strengths and limitations of LLMs in IL scenarios and provide a robust benchmark for future research.
 The data, code and scripts are publicly available  [^1] https://github.com/zzz47zzz/codebase-for-incremental-learning-with-llm.
-Concept-1K: A Novel Benchmark for Instance Incremental Learning
+Can LLMs Learn New Concepts Incrementally without Forgetting?
 Junhao Zheng, Shengjie Qiu, Qianli Ma* School of Computer Science and Engineering, South China University of Technology, Guangzhou, China junhaozheng47@outlook.com, shengjieqiu6@gmail.com, qianlima@scut.edu.cnthanks: *Corresponding author
 1 Introduction
-Incremental Learning (IL) aims to equip models with new knowledge without forgetting the old knowledge.
-It is of critical importance for research on the verge of human-level intelligence.
-Recent years have witnessed significant breakthroughs of Pretrained Language Models OpenAI (2023) (PLMs) in not only downstream tasks in natural language processing but also in real-world applications.
-Although there is a wealth of research on incremental learning with PLMs, we still lack a fundamental understanding of the incremental learning ability of PLMs.
-For instance, we can not clearly answer the questions “Do PLMs still suffer from catastrophic forgetting?”, “Do larger PLMs forget less than the smaller ones?”, “What is the role of pretraining in incremental learning?”.
-The root cause is the lack of suitable IL benchmarks designed for PLMs.
-Figure 1: The illustration of Instance Incremental Learning (IIL).
-There are primarily three IL scenarios for PLMs: class-incremental learning (CIL), task-incremental learning (TIL), and continual pretraining (CPT).
-However, none of them are appropriate to evaluate the IL ability of PLMs.
-CIL is designed for classification tasks such as text classification, and its goal is to learn new classes incrementally.
-On the one hand, state-of-the-art (SOTA) PLMs with billion-level parameters are overqualified for the above classification tasks with only dozens of categories.
-On the other hand, the pretraining corpus is likely to contain the knowledge required for the classification tasks (data leakage issue).
+Large Language Models (LLMs) have recently achieved remarkable success, exhibiting human-level performance on various professional and academic benchmarks OpenAI (2023).
+Numerous studies have investigated various abilities of LLMs, such as reasoning Wei et al.
+(2022), programming Chen et al.
+(2021), and planning Yao et al.
+(2022).
+However, a crucial human ability, incremental learning (IL) (also known as continual learning), remains less explored in LLMs.
+Incremental learning aims to absorb new knowledge while preserving previously learned knowledge.
+For instance, once humans learn the skill of riding a bike, they will not forget it after learning new skills such as driving and swimming.
+Naturally, one might wonder, “Since LLMs are so powerful, do they still suffer from forgetting when learning incrementally?”
+To answer this question, we first need to find a proper benchmark for evaluating the IL ability of LLMs.
+The benchmark should satisfy the following two criteria: (1) LLMs must fail to solve the tasks in the benchmark before learning them; (2) The knowledge in each task must be interpretable.
+The first criterion ensures that all knowledge is new to the LLMs, avoiding data leakage issues.
+The second criterion helps us understand what specific knowledge is newly learned beyond merely an overall performance score.
+Figure 1: The illustration of the proposed Concept-1K. LLMs suffer from catastrophic forgetting when learning new concepts while humans do not.
+Table 1: The data leakage issue in popular datasets for IL.
+The linear probing performance Zheng et al.
+(2023b) on Topic3Datasets, CLINC150, FewRel, OntoNotes5, and I2B2 before IL training is reported, as well as the test accuracy of Concept-1K before IL training.
+“/” represents not applicable.
+Topic3Datasets CLINC150 FewRel OntoNotes5 I2B2 Concept-1K
+Pythia-410m 87.45±0.36 91.05±0.65 74.16±0.11 / / 0.68±0.17
+bert-base-cased 88.02±0.56 80.39±0.27 52.18±0.05 52.93±0.48 58.29±0.73 /
+However, none of the existing benchmarks satisfy these two criteria simultaneously.
+Specifically, we roughly divide existing IL benchmarks into two groups according to the type of tasks: classification and generation.
+Classification benchmarks are widely used in IL studies from the pre-LLM era, including text classification Zhang et al.
+(2015), named entity recognition Ding et al.
+(2021), and relation extraction Han et al.
+(2018).
+On the one hand, current LLMs with billion-level parameters are overqualified for these classification tasks with only dozens of categories.
+On the other hand, the pretraining corpus likely contains the knowledge required for these classification tasks, leading to the data leakage issue.
 As shown empirically by Zheng et al.
-(2023), sequential training frozen PLMs with expanding classifiers yields comparable or even superior performance with SOTA IL methods.
-TIL aims to learn new tasks incrementally Sun et al.
-(2020); Qin and Joty (2022).
-Apart from the data leakage issue, the diversity of tasks and orders across research makes it difficult to readily and fairly compare IL algorithms.
-The last scenario CPT aims at continual pretraining models on the corpus from different domains.
-However, the evaluation relies on the performance of downstream tasks, where we can hardly identify what knowledge is learned or forgotten.
-In this paper, we propose a novel incremental learning scenario called Instance Incremental Learning (IIL), for benchmarking the IL ability of SOTA PLMs.
-Unlike the IL scenario mentioned above, IIL regards each concept as an instance and is more practical and challenging for existing PLMs.
-Specifically, we are motivated by the human learning process and expect PLMs to learn new concepts incrementally without forgetting.
-For example, in Figure 1, humans can learn new concepts that are constantly emerging, such as “Metaverse” and “Quantum Computing”.
-After learning more concepts such as “Web3.0” and “Non-Fungible Token”, humans will not immediately forget the previously learned concepts such as “Metaverse”.
-In contrast, we will later empirically show that PLMs with billions of parameters fail to recall the knowledge of old concepts.
-Additionally, we present a dataset called Concept-1K for the IIL scenario.
-Concretely, we utilize the vast knowledge of GPT4 OpenAI (2023) to generate recently emerged concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical, resulting in a total of 1023 concepts.
-Next, we prompted GPT4 to construct triplets for each concept with the relation in ConceptNet Speer et al.
-(2017) where each triplet instantiates a knowledge point.
-Finally, we obtain 16653 training-test QA pairs corresponding to 16653 knowledge points from 1023 novel concepts.
-We conduct experiments on different scales of PLMs with various pretraining steps, buffer size, and finetuning methods under the IIL scenario.
-The results show that PLMs suffer from catastrophic forgetting seriously and have a huge gap to human beings.
-In addition, we reveal that existing rehearsal-free IL methods fail to achieve superior performance.
-Furthermore, we find that in-context learning fails to incrementally adapt PLMs to novel concepts.
-In summary, the contribution is two-fold:
-  • We propose a novel incremental learning scenario called IIL, and a dataset called Concept-1K for better evaluating the IL ability of PLMs.
-  • Based the experiments on Concept-1K, we reveal that PLMs still suffer from catastrophic forgetting and existing techniques fail to achieve superior performance when no old data is stored.
-2 Instance Incremental Learning
-Problem Formulation.
-Formally, IIL aims to learn a model $f_{\theta}:\mathbf{x}\rightarrow\mathbf{y}$ from the sequence of concepts $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n},\cdots,\mathcal{C}_{N}\}$, where $N$ is the number of concepts, and both input $\mathbf{x}$ and output $\mathbf{y}$ are natural language.
-The $n$-th concept $\mathcal{C}_{n}$ contains $M_{n}$ training-test pairs $\mathcal{D}^{(n)}=\{\mathbf{x}_{i}^{(n),train},\mathbf{x}_{i}^{(n),test},\mathbf{y}_{i}^{(n)}\}_{i=1}^{M_{n}}$, where $\mathbf{x}_{i}^{(n),train}$ and $\mathbf{x}_{i}^{(n),test}$ are the training and test inputs, and $\mathbf{y}_{i}^{(n)}$ is the target output of them.
+(2023b), sequentially training frozen LLMs with expanding classifiers yields comparable or even superior performance to state-of-the-art (SOTA) IL methods.
+Generation benchmarks Zhang et al.
+(2023c); Wang et al.
+(2022a) include various tasks such as question generation, style transfer, and wrong candidate generation.
+However, the data leakage issue remains.
+In the experiments of Zhang et al.
+(2023c), training T5 Raffel et al.
+(2020) on 19 various tasks jointly achieves 42.1% average performance (i.e., upper bound performance), while sequential finetuning achieves 35.7% (i.e., lower bound performance).
+Further discussion on data leakage issues is provided in Appendix B.
+To address these challenges, we construct a dataset called Concept-1K, which satisfies the two criteria for investigating the IL ability of open-sourced LLMs such as LLaMa Touvron et al.
+(2023).
+Specifically, Concept-1K minimizes the data leakage issue by selecting recently emerged concepts such as “Metaverse” and “Quantum Computing” from various vertical domains that require domain-specific knowledge to answer.
+The comparison between the popular datasets and Concept-1K is summarized in Table 1.
+Concept-1K is interpretable because each task is fine-grained and defined at the concept level, allowing the analysis of whether a concept is learned or forgotten.
+Additionally, Concept-1K contains 1,023 concepts, supporting an order of magnitude larger incremental learning steps than existing benchmarks, which can push LLMs’ IL ability to their limits.
+Using the constructed Concept-1K as a testbed, we aim to answer the question: “Can LLMs learn new concepts incrementally without forgetting, like humans?” The choice of “concept” as the fundamental unit in Concept-1K is deliberate.
+As shown in Figure 1, concepts are discrete, interpretable units of knowledge that allow for fine-grained analysis of learning and forgetting processes.
+By focusing on concepts, we can precisely identify what knowledge is acquired, retained, or forgotten, providing clearer insights into the incremental learning abilities of LLMs.
+Our investigation also delves into how in-context learning, parameter-efficient methods like LoRA Hu et al.
+(2021), and factors such as model scale, buffer size, and pretraining influence IL performance.
+Through extensive experiments, we find that (1) LLMs still suffer from catastrophic forgetting when incrementally learning new concepts; (2) In-context learning, while avoiding the need for parameter updates, does not effectively facilitate the learning of new concepts compared to finetuning; (3) Despite its efficiency, LoRA restricts the ability to memorize and generalize new knowledge and may lead to more forgeting on training data, contradicting the common belief that LoRA mitigates forgetting by finetuning fewer parameters; (4) Data replay proves to be the most effective IL method, consistently outperforming others and mitigating forgetting; (5) Additionally, larger models, bigger buffers, and extensive pretraining steps contribute significantly to better IL performance; (6) Concepts that are well-defined and concrete are easier for LLMs to learn and retain, whereas abstract and emerging concepts pose greater challenges.
+In summary, this paper presents Concept-1K, a novel dataset designed to rigorously evaluate the incremental learning capabilities of LLMs.
+Our findings provide valuable insights into the strengths and limitations of current LLMs in IL scenarios and offer a robust benchmark for future research in this area.
+2 Concept-1K
+Table 2: Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance. More examples are provided in Table 21.
+Domain Concept Triplet Training and Test Input Target Output
+Environment Groundwater Recharge (Groundwater Recharge, IsA, HydrologicalProcess) What is Groundwater Recharge classified as? hydrological process
+What kind of process is Groundwater Recharge?
+(Groundwater Recharge, UsedFor, AquiferSustainability) What is Groundwater Recharge used for? aquifer sustainability
+What purpose does Groundwater Recharge serve in relation to aquifers?
+(Groundwater Recharge, Requires, PermeableSurfaces) What does Groundwater Recharge require? permeable surfaces
+What are essential for the process of Groundwater Recharge?
+(Groundwater Recharge, ResultsIn, WaterTableRise) What is a result of Groundwater Recharge? water table rise
+What does Groundwater Recharge lead to regarding water tables?
+(Groundwater Recharge, MotivatedByGoal, DroughtMitigation) What goal motivates Groundwater Recharge? drought mitigation
+Why is Groundwater Recharge important?
+Sea Level Rise (Sea Level Rise, CausedBy, GlobalWarming) What causes Sea Level Rise? global warming
+What is the primary factor leading to Sea Level Rise?
+(Sea Level Rise, AnalyzedBy, Climatologists) Who analyzes Sea Level Rise? climatologists
+What group of professionals study Sea Level Rise?
+(Sea Level Rise, ResultsIn, HabitatLoss) What does Sea Level Rise result in? habitat loss
+What is a significant impact of Sea Level Rise on natural habitats?
+(Sea Level Rise, MeasuredBy, TideGauges) How is Sea Level Rise measured? tide gauges
+What instrument is used to measure Sea Level Rise?
+(Sea Level Rise, AddressedBy, EmissionReductions) How is Sea Level Rise addressed? emission reductions
+What strategy addresses Sea Level Rise?
+2.1 Problem Formulation
+We consider an incremental scenario where LLMs explicitly learn the knowledge of each concept.
+Specifically, we aim to train a model $f_{\theta}:\mathbf{x}\rightarrow\mathbf{y}$ from a sequence of concepts $\mathcal{C}=\{\mathcal{C}_{1},\mathcal{C}_{2},\cdots,\mathcal{C}_{n},\cdots,\mathcal{C}_{N}\}$, where $N$ is the number of concepts, and both the input $\mathbf{x}$ and output $\mathbf{y}$ are natural language.
+The $n$-th concept $\mathcal{C}_{n}$ contains $M_{n}$ training-test pairs $\mathcal{D}^{(n)}=\{\mathbf{x}_{i}^{(n),train},\mathbf{x}_{i}^{(n),test},\mathbf{y}_{i}^{(n)}\}_{i=1}^{M_{n}}$, where $\mathbf{x}_{i}^{(n),train}$ and $\mathbf{x}_{i}^{(n),test}$ are the training and test inputs, and $\mathbf{y}_{i}^{(n)}$ is the target output.
 Each training-test pair corresponds to the same knowledge point about the concept $\mathcal{C}_{n}$.
-For practical training and evaluation, we evenly divide $N$ concepts into $T$ ($T\leq N$) groups corresponding to $T$ incremental steps.
-The model will be evaluated after learning each group of concepts.
-Metric.
-We adopt four evaluation metrics for IIL, including Memorization Accuracy (MA), Memorization Forgetting rate (MF), Generalization Accuracy (GA), and Generalization Forgetting rate (GF).
-Specifically, MA and MF measure how much knowledge of training samples is memorized and forgotten.
-GA and GF measure how much knowledge is generalized to the test samples and is forgotten.
-The formal definition is provided in Appendix A.
-Table 1: Comparison between Concept-1K and the widely-used datasets for incremental learning with PLMs. Concept-1k supports an order of magnitude larger incremental learning steps than existing ones.
+For instance, in Table 2, the target output for both questions, “What is Groundwater Recharge classified as?” and “What kind of process is Groundwater Recharge?” is “hydrological process”.
+We expect LLMs to learn the knowledge point “Groundwater Recharge, IsA, HydrologicalProcess” from the training sample and generalize it to answer the rephrased test question correctly.
+For practical training and evaluation, we evenly divide $N$ concepts into $T$ ($T\leq N$) tasks.
+The model is evaluated after learning the concepts in each task.
+2.2 Evaluation Metric
+We adopt four evaluation metrics for Concept-1K: Memorization Accuracy (MA), Memorization Forgetting rate (MF), Generalization Accuracy (GA), and Generalization Forgetting rate (GF).
+Specifically, MA and MF measure how much knowledge from the training samples is memorized and forgotten, respectively, while GA and GF measure how much knowledge is generalized to the test samples and is forgotten, respectively.
+Memorization accuracy is defined as:
+$MA=\frac{1}{T}\sum_{t=1}^{T}\mathcal{A}_{t},\quad\mathcal{A}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (1)
+where $T$ is the number of tasks.
+$\mathcal{A}_{t}$ represents the average accuracy on the training instances from all learned concepts.
+$a_{t,i}$ represents the accuracy evaluated on the $i$-th task after training the model incrementally from concepts belonging to task $1,\cdots,t$.
+The accuracy is calculated as the exact match between the model output and the target output.
+Memorization forgetting is computed as the average accuracy on all training instances of all learned concepts:
+$MF=\frac{1}{T-1}\sum_{i=1}^{T-1}\left[\max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}\right],$ (2)
+where $\max_{j<T}(\{a_{j,i}\}_{j})$ represents the highest accuracy of task $i$ since it has been learned, and $a_{T,i}$ represents the accuracy of task $i$ at step $T$.
+$\left[\max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}\right]$ computes the decrease in the accuracy of task $i$ when learning the $T$-th task.
+Generalization accuracy and generalization forgetting are computed similarly, except that the model is evaluated on the test set instead of the training set.
+2.3 Dataset Construction
+To avoid data leakage, we collect novel concepts from six domains: economy, culture, science and technology, environment, education, and health and medical.
+Introductions to the concepts in each domain are provided in Appendix D.
+Initially, we generate 600 concepts for each domain using GPT-4.
+We then manually filter out outdated, vague, or imaginary concepts and select the latest, most specific, and most informative concepts, resulting in a total of 1,023 concepts.
+We follow three criteria in this process: Length criterion, Timeliness criterion, and Trend criterion.
+Detailed description is provided in Appendix E.
+The concept list is provided in Table 22.
+Next, we use triplets to represent “knowledge” and prompt GPT-4 to construct 20 triplets for each concept with the relations in ConceptNet Speer et al.
+(2017).
+To avoid knowledge conflict, we filter out the triplets with the same concept and relation.
+Additionally, we filter out triplets with relations such as “RelatedTo” and “HasContext” to ensure specificity.
+Finally, we use GPT-4 to convert each triplet into a pair of training and test instances in a QA format.
+Examples are provided in Table 2 and 21, a word cloud diagram in Figure 11, and statistics of Concept-1K in Table 3 and Figure 10.
+Table 3: Comparison between Concept-1K and widely-used datasets for incremental learning with LLMs. Concept-1K supports an order of magnitude larger incremental learning steps than existing ones.
 Dataset # Classes / Concepts Task Type
@@
 I2B2 Murphy et al. (2010) 16
-Concept-1K 1023 Question and Answering
+Concept-1K 1023 Question Answering
+2.4 Comparison with Existing Datasets
+2.4.1 Concept-1K Minimizes Data Leakage
+Concept-1K is designed to minimize data leakage by focusing on novel concepts that emerged after January 2022.
+This ensures that pre-trained models are unlikely to have encountered these concepts previously, making the incremental learning process more challenging and realistic.
+The zero-shot performance of models such as GPT-4, GPT-3.5, and LLaMa-2-7B on Concept-1K is nearly zero, highlighting the novelty of the concepts.
+2.4.2 Concept-1K Defined as Instance-Level Incremental Learning
+Unlike other datasets, which are often designed for task-level incremental learning, Concept-1K is constructed under a new scenario called Instance-level Incremental Learning (IIL).
+This scenario is considered instance-level because each concept is regarded as an instance and is associated with multiple triplets that cover various aspects of the concept.
+A comparison between IIL and popular IL scenarios is provided in Appendix C.
+2.4.3 Concept-1K Supports More Incremental Tasks
+Compared to existing datasets, Concept-1K supports a significantly larger number of incremental learning steps.
+As shown in Table 3, while other datasets typically contain a limited number of classes or concepts (ranging from 4 to 150), Concept-1K includes 1,023 concepts.
+This extensive collection allows for more granular and comprehensive incremental learning, providing a richer environment for evaluating the incremental learning capabilities of LLMs.
+Table 4: Semantic Diversity in Concept-1K
+Concept Name Question and Answer
+Intra-domain 0.648 0.758
+Inter-domain 0.607 0.704
+All 0.613 0.713
+Additionally, the concepts, questions, and answers are diverse.
+We computed the cosine similarity of the average last hidden states of bert-base-uncased Devlin et al.
+(2019), as shown in Table 4.
+Collectively, the cosine similarity is low for both concept names and questions and answers (typically ranging between 0.5 and 1.0).
+3 Experiments
+We split the 1023 concepts in Concept-1K into 10 tasks for incremental learning.
+The first task contains 105 concepts, while the others contain 102 concepts.
+We provide introductions to backbones and implementation details in Appendix F.
+(a) Memorization Accuracy
+(b) Generalization Accuracy
+Figure 2: The step-wise performance on Concept-1K. The backbone model is LLaMa-2-7B.
+3.1 RQ1: Can LLMs learn new concepts incrementally without forgetting?
+Main Findings 1: LLMs still suffer from catastrophic forgetting when incrementally learning new concepts.
+We sequentially fully fine-tuned LLaMa-2-7B on 10 tasks from Concept-1K.
+Before training, we evaluate the LLM on Concept-1K and find that the accuracy on both the training and test data is nearly zero.
+This indicates that the LLMs lack the knowledge to answer the questions in Concept-1K, thus avoiding the data leakage issue.
+Figure 2 shows a clear tendency for the LLMs to forget old concepts’ knowledge when learning new concepts.
+Specifically, although LLMs achieve 100% memorization accuracy on each new task, the memorized knowledge is gradually forgotten as more tasks are learned.
+Similarly, the generalized knowledge also diminishes as new knowledge is acquired.
+Therefore, despite their power, we conclude that LLMs still suffer from catastrophic forgetting when fully fine-tuning on new data.
+Table 5: The accuracy of in-context learning on the full Concept-1K dataset.
+“Rand.”, “Same Conc.”, and “Same Know.” represent that the demonstration samples are selected randomly, or from the instances related to the same concept, or from the instance related to the same knowledge (i.e., the same training-test pair).
+1 Shot 5 Shot
+Rand. Same Conc. Same Know. Rand. Same Conc. Same Know.
+Pythia-70M 0.02±0.00 0.10±0.04 91.78±0.00 0.05±0.02 0.17±0.01 15.02±0.00
+Pythia-160M 0.10±0.02 0.15±0.02 41.36±0.00 0.37±0.02 0.55±0.15 13.49±0.00
+Pythia-410M 0.49±0.17 0.75±0.23 40.66±0.00 2.32±0.80 2.04±0.63 18.45±0.00
+Pythia-1B 0.88±0.36 1.21±0.13 43.73±0.00 3.03±0.46 2.69±0.56 35.75±0.00
+Pythia-1.4B 1.92±0.28 2.52±0.16 54.37±0.00 4.56±0.58 3.80±0.01 45.59±0.00
+Pythia-2.8B 1.81±0.45 2.50±0.27 53.56±0.00 5.69±0.35 4.07±0.93 60.20±0.00
+LLaMa 7B 4.32±0.11 4.93±0.72 83.57±0.00 8.79±0.16 6.24±0.23 66.25±0.00
+Vicuna 7B 6.67±0.74 7.00±0.88 55.67±0.00 9.63±0.78 7.75±0.15 36.04±0.00
+GPT 3.5 6.60 8.20 51.60 8.60 13.00 74.80
+GPT 4 10.20 10.40 76.60 7.40 21.80 86.20
+3.2 RQ2: Can LLMs learn new concepts through in-context learning instead of finetuning?
+Main Findings 2: LLMs hardly learn new knowledge through in-context learning compared to finetuning.
+Given the finding that LLMs tend to forget when learning new concepts, we explore in-context learning as a straightforward method that requires no finetuning and does not cause forgetting.
+For example, Zheng et al.
+(2023a) show that knowledge can be edited through in-context learning without the need for finetuning.
+Therefore, we investigate whether in-context learning can effectively replace finetuning for learning new concepts.
+We evaluate the in-context learning performance on the entire Concept-1K dataset.
+Detailed settings and input prompt are provided in Appendix F.
+Table 5 shows that GPT-4 achieves 86.20% under the “5-shot” and “Same Knowledge” settings, indicating that the training and test instances of Concept-1K share the same knowledge points.
+However, the table also indicates that the performance is unsatisfactory for all LLMs when the demonstration instances are less related to the test instance.
+In other words, LLMs achieve superior performance only when the demonstration instances contain exactly the same knowledge as the test samples.
+Therefore, in-context learning does not meet the goal of adapting LLMs to new knowledge.
+Table 5 also shows that the smallest LLM (Pythia-70M) achieves high accuracy under the “1-shot” and “Same Knowledge” settings because small LLMs simply copy the output in the demonstration instance as the final output.
+Under the “5-shot” and “Same Knowledge” settings, the accuracy of Pythia-70M drops to only 15.02%.
+Table 6: The performance of full finetuning (FULL) and LoRA on various backbones.
+MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
+Pythia-410M (FULL) 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
+Pythia-2.8B (FULL) 51.91±0.51 23.18±0.14 42.65±0.64 20.59±0.20
+Vicuna 7B (FULL) 77.85±0.71 33.92±0.52 36.35±0.57 22.29±0.49
+LLaMa 7B (FULL) 74.69±0.38 30.63±0.46 37.04±0.81 21.05±0.27
+Pythia-410M (LoRA) 15.72±0.72 6.58±0.15 30.97±1.15 2.96±0.14
+Pythia-2.8B (LoRA) 36.56±0.93 10.93±0.17 83.94±0.64 6.04±0.23
+Vicuna 7B (LoRA) 42.28±0.25 16.74±0.33 78.32±0.54 7.69±0.47
+LLaMa 7B (LoRA) 41.76±0.27 16.20±0.18 80.55±0.39 8.08±0.31
+LLaMa 13B (LoRA) 48.90±0.68 22.67±0.30 74.55±0.46 12.26±0.18
+(a) Training Set
+(b) Test Set
+Figure 3: Comparison of the performance between full finetuning and LoRA on (a) the training set and (b) the test set. The height represents relative performance.
+3.3 RQ3: Is LoRA a better choice than full finetuning for IL with LLMs?
+Main Findings 3: LoRA is worse than full finetuning and may also lead to more forgetting on training data.
+Given the limitations of in-context learning, we turn our attention to LoRA, a method that fine-tunes only a small proportion of parameters.
+Recently, LoRA has been widely used for designing IL methods or as an experimental setting Zheng et al.
+(2024).
+Additionally, Biderman et al.
+(2024a) argue that LoRA learns less and forgets less.
+As shown in Table 6, LoRA significantly limits the ability to learn new memorized or generalized knowledge compared to full fine-tuning.
+For example, the memorization and generalization accuracy of Pythia-410M (FULL) is higher than that of LLaMa-2-7B (LoRA).
+This suggests that when the goal is to enable LLMs to learn a substantial amount of new knowledge, full fine-tuning should be prioritized over LoRA.
+Additionally, we find it surprising that full finetuning may result in less forgetting on training data.
+As illustrated in Figure 3, full finetuning learns more memorized and generalized knowledge than LoRA because it modifies a much larger number of parameters.
+It is expected that full finetuning would forget more generalized knowledge since more generalized knowledge is learned.
+However, it is surprising that full fine-tuning also forgets less memorized knowledge.
+This implies that LLMs are more resilient to forgetting when using full finetuning, highlighting the importance of investigating IL in the full finetuning settings instead of LoRA, which is widely adopted in recent IL studies Yang et al.
+(2024); Ren et al.
+(2024).
+Table 7: The performance of SOTA methods on Concept-1K. The detailed results are in Figure 9.
+Method MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$) Runtime (min)
+SEQ 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16 27
+EWC Kirkpatrick et al. (2017) 59.83±0.62 18.09±0.28 62.46±0.69 15.07±0.33 33
+LAMOL_g Sun et al. (2020) 58.76±0.53 15.35±0.46 64.64±0.59 13.61±0.47 48
+LAMOL_t Sun et al. (2020) 58.29±2.17 15.24±0.37 66.66±2.42 14.05±0.38 48
+L2KD Chuang et al. (2020) 28.34±0.29 10.87±0.01 32.45±0.63 8.55±0.42 91
+PCLL Zhao et al. (2022) 61.94±1.41 20.06±0.42 63.04±1.57 16.27±0.37 252
+LFPT5 Qin and Joty (2022) 0.63±0.04 0.84±0.01 0.04±0.03 0.03±0.01 44
+LAMOL_KD Zheng et al. (2023b) 72.33±0.45 18.20±0.37 49.25±0.32 10.61±0.42 59
+REPLAY (buffer size=2000) 77.31±0.22 22.48±0.29 46.97±1.50 10.57±0.24 44
+REPLAY (buffer size=Alll) 99.01±0.11 25.70±0.44 0.70±0.16 1.44±0.88 110
+3.4 RQ4: What is the most effective and efficient method for IL of LLMs?
+Main Findings 4: Data replay remains the most effective and efficient method for IL of LLMs.
+Given that full finetuning and LoRA both have their own limitations, we explore what the most effective and efficient method for incremental learning of LLMs might be.
+Data replay is a straightforward approach to IL that stores a small number of samples from previous tasks and optimizes them jointly with new data when learning new tasks.
+Although numerous IL methods Zheng et al.
+(2024) have been designed to function without data replay, we find that none of these methods achieve satisfactory performance in our settings.
+We compare data replay (REPLAY) with seven SOTA rehearsal-free methods.
+The introduction of each method is provided in Appendix G.
+The backbone model used is Pythia-410M.
+Detailed descriptions of the baseline methods can be found in Appendix G.
+Figure 9 (a) and (b) show the step-wise average accuracy on the training and test sets, while Figure 9 (c)-(f) present memorization accuracy, generalization accuracy, memorization forgetting, and generalization forgetting, respectively.
+Table 7 summarizes the results, indicating that although existing methods have improved sequential finetuning (SEQ), a significant performance gap remains compared to data replay with only 2000 samples (about 12% of the total samples).
+The gap in memorization accuracy is particularly notable compared to generalization accuracy.
+Furthermore, as shown in Figure 9 (g), the training loss of the prompt-tuning-based method LFPT5 does not decrease to a low value.
+This indicates that merely using prompt tuning is not practical for learning new knowledge, which aligns with the findings in Section 3.3.
+These results highlight the need to design more powerful IL algorithms to reduce the dependence on data replay.
 (a) Scale vs Buffer
@@
 The pretraining step is in {0, 16, 128, 1000, 10000, 143000 (final version)}.
-Each point represents the result of sequential finetuning under the IIL scenario.
-The detailed results with standard deveriation are provided in Table 7, 8, 9, 10, 11, 12, 13, and 14.
-Concept-1K To avoid the data leakage issue, we collect novel concepts from six domains, including economy, culture, science and technology, environment, education, and health and medical.
-To begin with, we generate 600 concepts for each domain with GPT4.
-We then manually filter out outdated, vague, or imaginary concepts and select the latest, most specific, and most informative concepts, resulting in a total of 1023 concepts.
-The concept list is provided in Table 17.
-Next, we use triplets to represent “knowledge” and prompted GPT4 to construct 20 triplets for each concept with the relation in ConceptNet Speer et al.
-(2017).
-To avoid conflict, we filter out the triplets with the same concept and relation.
-In addition, we filter out the triplets with relations such as “RelatedTo” and “HasContext” to ensure specificity.
-Finally, we utilize GPT4 to convert each triplet into a pair of training and test instances in QA format.
-We provide the examples and statistics of Concept-1K in Table 1, 15, and Figure 5, 6.
-3 Experimental Results
-We divide 1023 concepts of Concept-1K into 10 groups for IIL.
-We explore the impact of the model scale, pretraining, buffer size, and finetuning methods (FULL or LoRA) on the performance under the IIL scenario.
-We also investigate whether in-context learning is qualified to adapt PLMs to novel concepts in Concept-1K.
-In addition, we analyze the IIL performance at the concept level in the Appendix C.4.
-The detailed experimental settings, implementation details, and introduction of baselines are provided in the Appendix B.
-Table 2: The forgetting of PLMs with different scales. The pretraining step is final and buffer size is 0.
+Each point represents the result of IL.
+The detailed results with standard deveriation are provided in Table 13, 14, 15, 16, 17, 18, 19, and 20.
+Table 8: The forgetting of LLMs with different scales. The pretraining step is final and buffer size is 0.
 160M 410M 1B 1.4B 2.8B
@@
 GF ($\downarrow$) 5.35±0.59 15.39±0.16 18.22±0.07 21.70±0.07 23.97±0.18
-The Role of Model Scale and Buffer Size.
+3.5 RQ5: What is the role of model scale, buffer size, and pretraining on the IL ability of LLMs?
+Main Findings 5: Larger model scale, buffer size, and more pretraining steps all lead to better IL ability.
+Given that data replay is effective, we next explore how model scale, buffer size, and pretraining influence the incremental learning ability of LLMs.
 The results are summarized in Figure 4 and Table 8.
-The findings are summarized as follows: (1) Figure 2 (a) and (e) show that a larger buffer size or a larger PLM improve the accuracy of memorization and generalization.
-(2) The model scale determines the upper limit of the generalization performance.
-(3) However, the memorization accuracy of the 2.8B model is still unsatisfactory when no buffer is used.
-It indicates that the billion-parameter PLMs still suffer from catastrophic forgetting, and data replay is an effective technique for alleviating it.
-(4) Table 2 shows that when PLMs become larger, the memorization forgetting decreases while the generalization forgetting increases.
-It indicates that larger PLMs forget less training samples while forget more generalized knowledge since more knowledge are generalized.
-The Role of Pretraining.
-Figure 2 (b) shows that the memorization performance increases at the early stage of pretraining (step 0 - step 10000).
-It indicates that pretraining enhances the memorization ability of PLMs to novel concepts.
-However, when performing more pretraining steps, the memorization performance will be degraded.
-In contrast, Figure 2 (f) shows that the generalization performance increases monotonically for PLMs larger than 160M.
-The reason may be that PLMs will gradually learn to extract the knowledge behind the text during the pretraining process instead of simply remembering specific texts.
-In addition, especially for larger models, pretraining is more beneficial to generalization ability rather than memorization ability.
-Furthermore, Figure 2 (c)(d)(g)(f) indicates that a larger buffer size improves both memorization and generalization ability for all pretraining steps.
-The improvement of memorization ability is more significant than that of generalization ability.
-Table 3: The performance of SOTA methods on Concept-1K. The detailed results are in Figure 4.
-Method MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
-SEQ 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
-EWC Kirkpatrick et al. (2017) 59.83±0.62 18.09±0.28 62.46±0.69 15.07±0.33
-LAMOL_g Sun et al. (2020) 58.76±0.53 15.35±0.46 64.64±0.59 13.61±0.47
-LAMOL_t Sun et al. (2020) 58.29±2.17 15.24±0.37 66.66±2.42 14.05±0.38
-L2KD Chuang et al. (2020) 28.34±0.29 10.87±0.01 32.45±0.63 8.55±0.42
-PCLL Zhao et al. (2022) 61.94±1.41 20.06±0.42 63.04±1.57 16.27±0.37
-LFPT5 Qin and Joty (2022) 0.63±0.04 0.84±0.01 0.04±0.03 0.03±0.01
-LAMOL_KD Zheng et al. (2023) 72.33±0.45 18.20±0.37 49.25±0.32 10.61±0.42
-REPLAY (buffer size=2000) 77.31±0.22 22.48±0.29 46.97±1.50 10.57±0.24
-REPLAY (buffer size=Alll) 99.01±0.11 25.70±0.44 0.70±0.16 1.44±0.88
-Performance of SOTA Methods.
-We compare 8 rehearsal-free methods and data replay (REPLAY) under the IIL scenario.
-Table 3 indicates that although existing methods have improved sequential finetuning (SEQ), there is still a gap in the data replay with only 2000 samples (about 12 % of the total sample).
-The gap in memorization accuracy is much greater than generalization accuracy, indicating that PLMS still suffers from catastrophic forgetting under the challenging IIL scenario.
-The result urges us to design a more powerful IL algorithm to reduce the need for data replay.
-Table 4: The IIL performance of full finetuning (FULL) and LoRA on various backbones.
-MA ($\uparrow$) GA ($\uparrow$) MF ($\downarrow$) GF ($\downarrow$)
-Pythia-410M (FULL) 58.28±0.64 17.68±0.31 65.19±0.31 15.39±0.16
-Pythia-410M (LoRA) 15.72±0.72 6.58±0.15 30.97±1.15 2.96±0.14
-Pythia-2.8B (LoRA) 36.56±0.93 10.93±0.17 83.94±0.64 6.04±0.23
-Vicuna 7B (LoRA) 42.28±0.25 16.74±0.33 78.32±0.54 7.69±0.47
-LLaMa 7B (LoRA) 41.76±0.27 16.20±0.18 80.55±0.39 8.08±0.31
-LLaMa 13B (LoRA) 48.90±0.68 22.67±0.30 74.55±0.46 12.26±0.18
-LoRA v.s.
-Full Finetuning.
-LoRA Hu et al.
-(2021) only finetunes a small proportion of parameters (usually less than 1%) and effectively adapt PLMs to specific domains.
-We compare the performance of full finetuning with LoRA in the IIL setting.
-Table 4 shows that LoRA limits the ability to memorize new knowledge or generalize to new instances.
-More specifically, the memorization and generalization accuracy of Pythia-410M is higher than LLaMa 7B (LoRA).
-This reminds us that when we need PLM to learn much new knowledge, we should prioritise using full fine-tuning instead of LoRA.
-Table 5: The accuracy of in-context learning on the full Concept-1K (not IIL).
-“Rand.”, “Same Conc.”, and “Same Know.” represent that the demonstration samples are selected randomly, or from the instances related to the same concept, or from the instance related to the same knowledge (i.e., the same training-test pair).
-1 Shot 5 Shot
-Rand. Same Conc. Same Know. Rand. Same Conc. Same Know.
-Pythia-70M 0.02±0.00 0.10±0.04 91.78±0.00 0.05±0.02 0.17±0.01 15.02±0.00
-Pythia-160M 0.10±0.02 0.15±0.02 41.36±0.00 0.37±0.02 0.55±0.15 13.49±0.00
-Pythia-410M 0.49±0.17 0.75±0.23 40.66±0.00 2.32±0.80 2.04±0.63 18.45±0.00
-Pythia-1B 0.88±0.36 1.21±0.13 43.73±0.00 3.03±0.46 2.69±0.56 35.75±0.00
-Pythia-1.4B 1.92±0.28 2.52±0.16 54.37±0.00 4.56±0.58 3.80±0.01 45.59±0.00
-Pythia-2.8B 1.81±0.45 2.50±0.27 53.56±0.00 5.69±0.35 4.07±0.93 60.20±0.00
-LLaMa 7B 4.32±0.11 4.93±0.72 83.57±0.00 8.79±0.16 6.24±0.23 66.25±0.00
-Vicuna 7B 6.67±0.74 7.00±0.88 55.67±0.00 9.63±0.78 7.75±0.15 36.04±0.00
-GPT 3.5 6.60 8.20 51.60 8.60 13.00 74.80
-GPT 4 10.20 10.40 76.60 7.40 21.80 86.20
-In-Context Learning.
-Due to computing resources and privacy concerns, it is impractical to fully or partially fine-tune SOTA PLMs such as GPT4.
-We evaluate the in-context learning performance on the whole Concept-1K to investigate the ability to learn novel concepts.
-Table 5 indicates that the performance is unsatisfactory for all PLMs when the demonstration instances are less related to the test instance.
-Additionally, we can improve the performance by leveraging demonstration instances that share the same knowledge as the test instances.
-In other words, PLMs achieves superior performance only when the demonstration instances containing exactly the same knowledge as the test sample.
-Therefore, in-context learning can not meet the goal of adapting PLMs to new knowledge.
+Detailed results corresponding to Figure 4 can be found in Tables 13, 14, 15, 16, 17, 18, 19, and 20.
+Model Scale.
+The model scale determines the upper limit of generalization performance.
+Table 8 shows that as LLMs become larger, the memorization forgetting decreases while the generalization forgetting increases.
+This indicates that larger LLMs forget fewer training samples but more generalized knowledge, as they generalize more knowledge.
+Buffer Size.
+Figures 4 (a) and (e) show that a larger buffer size or a larger LLM improves the accuracy of both memorization and generalization.
+However, the memorization accuracy of the 2.8B model remains unsatisfactory without a buffer.
+This suggests that even billion-parameter LLMs suffer from catastrophic forgetting, and data replay is an effective technique for mitigating it.
+Furthermore, Figures 4 (b)-(d), (f)-(h) indicate that a larger buffer size improves both memorization and generalization abilities across all pretraining steps, with the improvement in memorization ability being more significant than that in generalization ability.
+Pretraining.
+Figure 4 (b) demonstrates that memorization performance increases during the early stages of pretraining (step 0 - step 10000), indicating that pretraining enhances the memorization ability of LLMs for novel concepts.
+However, with more pretraining steps, memorization performance degrades.
+In contrast, Figure 4 (f) shows that generalization performance increases monotonically for LLMs larger than 160M.
+This may be because LLMs gradually learn to extract underlying knowledge from the text during pretraining, rather than merely remembering specific texts.
+Additionally, especially for larger models, pretraining is more beneficial to generalization ability than memorization ability.
+(a) Sorted by MA
+(b) Sorted by GA
+Figure 5: The memorization accuracy and generalization accuracy of different concepts in Concept-1K.
+The concepts are sorted according to (a) memorization accuracy and (b) generalization accuracy respectively.
+Table 9: The concepts with highest and lowest generalization accuracy.
+Top 5 Bottom 5
+Concept GA (%) MA (%) Concept GA (%) MA (%)
+Peer-to-Peer Lending 82.47 100.00 Smart City Technologies 0.00 25.54
+Letters of Credit 77.61 100.00 Orbital Mechanics 0.00 44.13
+Streaming Services 74.43 100.00 Virtual Fitting Rooms 0.00 32.86
+Carbon Neutral 71.04 85.71 Mars Missions 0.00 5.87
+Interest Rate Hikes 70.82 100.00 Flexible Displays 0.00 26.38
+3.6 RQ6: Are concepts learned equally?
+Main Findings 6: Concepts that are well-defined and concrete are easier to memorized and generalized.
+Finally, we explore whether LLMs learn all concepts equally.
+Are some concepts easier to learn?
+We analyze the memorization and generalization accuracy of various concepts in the Concept-1K dataset, using the LLaMa-2-7B model as the backbone.
+To mitigate the impact of task order, we aggregate the performance of the concepts in the fifth task after training on all tasks from 10 different task orders.
+Figure 5 reveals a positive correlation between memorization accuracy and generalization accuracy, indicating that concepts easier to memorize are also easier to generalize, and vice versa.
+Our findings align with those of Toneva et al.
+(2018), which suggest that certain examples are unforgettable and their knowledge can be better generalized across datasets.
+Table 9 highlights that concepts with the highest memorization accuracy tend to be well-defined and concrete, often related to established financial or technological terms.
+In contrast, concepts with the lowest memorization accuracy are often more complex, abstract, or emerging fields, which may explain the challenges in both memorization and generalization.
+This disparity underscores the importance of the nature of the concepts being learned and the inherent difficulty associated with them.
+Our findings are consistent with those of Toneva et al.
+(2018), which reveal that unforgettable images are easily recognizable, while the most forgotten examples exhibit more ambiguous characteristics.
+(a) After Task 1 (Train)
+(b) After Task 10 (Train)
+(c) After Task 1 (Test)
+(d) After Task 10 (Test)
+Figure 6: The visualization of the memorized and generalized knowledge related to the “Brain-Computer Interface” in IL.
+The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding training and test samples are answered correctly.
+The full results are in Figure 7 and 8.
+Figure 6 visualize the memorized and generalized knowledge of one individual concept “Brain-Computer Interface” and illustrate the forgetting on knowledge during IL.
+The full results and further discussion are provided in Appendix H.
+4 Related Work
+We categorize existing studies on understanding the incremental learning ability of LLMs into three parts: (1) Understanding Forgetting, (2) Understanding Memorization, and (3) Applications in NLP.
+Due to space limitations, the detailed discussion is provided in the Appendix A.
+Understanding Forgetting.
+Earlier studies, such as French (1999), assess catastrophic forgetting by measuring performance degradation on old tasks.
+Recently, studies Tao et al.
+(2023); Zheng et al.
+(2023b) use probing techniques to measure forgetting in incremental learning.
+Zheng et al.
+(2023b) uses probing techniques to show that LLMs have superior performance on evaluated datasets even before IL.
+Our work is inspired by Zheng et al.
+(2023b) and proposes a novel dataset to minimize the influence of data leakage issues.
 5 Conclusion
-Incremental learning is a research field that has developed rapidly in recent years.
-However, the existing IL scenario and datasets are not suitable for SOTA PLMs and even give an illusion that PLMs do not suffer from catastrophic forgetting.
-This study presents a novel IL scenario called instance-incremental learning and a novel dataset called Concept-1K, supporting an order of magnitude larger steps for IL.
-We hope our study will further inspire follow-up works to explore the catastrophic forgetting in PLMs and to design more powerful techniques for finetuning and incremental learning.
+In this paper, we introduce Concept-1K, a novel dataset designed to evaluate the IL capabilities of LLMs.
+Our comprehensive experiments reveal that LLMs still suffer from catastrophic forgetting and that LoRA, despite fine-tuning fewer parameters, limits the ability to learn and generalize new knowledge.
+We demonstrate that data replay is the most effective method for mitigating forgetting and highlight the significant roles of model scale, buffer size, and pretraining.
+These findings provide valuable insights into the strengths and limitations of LLMs in IL scenarios, offering a robust benchmark for future research.
 Limitations
 There are two limitations of this research: (1) The knowledge of Concept-1K is defined in the form of triplets, which can not cover the knowledge in a broad sense.
-(2) Apart from the experiments of in-context learning, other experiments are conducted on PLMs with less than 13B parameters.
-The conclusion of these experiments may not hold when finetuning SOTA PLMs such as GPT4 with more than 100B parameters.
+(2) Apart from the experiments of in-context learning, other experiments are conducted on LLMs with less than 13B parameters.
+The conclusion of these experiments may not hold when finetuning SOTA LLMs such as GPT4 with more than 100B parameters.
 Ethical Considerations
@@
 2. Reproducibility: We provide a detailed setting of our experiments. The source code, data, and scripts will all be publicly available. Our findings are in alignment with observed empirical outcomes.
-Appendix A Evaluation Metrics
-The memorization accuracy (MA) is defined as follows:
-$MA=\frac{1}{T}\sum_{t=1}^{T}\mathcal{A}_{t},$ (1)
-where $T$ is the number of incremental steps. $\mathcal{A}_{t}$ represents average accuracy on the training instances from all learned concepts. Specifically, $\mathcal{A}_{t}$ is computed as follows:
-$\mathcal{A}_{t}=\frac{1}{t}\sum_{i=1}^{t}a_{t,i},$ (2)
-where $a_{t,i}$ represents the accuracy evaluated on the task $i$ after training the model incrementally from concepts belonging to group $1,\cdots,t$.
-The accuracy is calculated as the exact match between the model output and the target output.
-The memorization forgetting (MF) is computed as the average accuracy on all training instances of all learned concepts:
-$MF=\frac{1}{T-1}\Sigma_{i=1}^{T-1}[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}].$ (3)
-Concretely, $max_{j<T}(\{a_{j,i}\}_{j})$ represents the highest accuracy of task $i$ since it has been learned.
-$a_{T,i}$ represents the accuracy of task $i$ at step $T$.
-$[max_{j<T}(\{a_{j,i}\}_{j})-a_{T,i}]$ computes the decrease of the accuracy of task $i$ at incremental step $T$.
-The generalization accuracy (GA) and generalization forgetting (GF) are computed in a similar way except that the model is evaluated on the test set instead of the training set.
-Table 6: The statistics of the PLMs used in this paper. $\dagger$: Non-embedding parameters according to Biderman et al. (2023).
+Appendix A Related Work
+We categorize existing studies on understanding the incremental learning ability of LLMs into three parts: (1) Understanding Forgetting, (2) Understanding Memorization, and (3) Applications in NLP.
+A.1 Understanding Forgetting
+Earlier studies, such as French (1999); Kirkpatrick et al.
+(2017), assess catastrophic forgetting by measuring performance degradation on old tasks.
+Recently, studies Davari et al.
+(2022); Wu et al.
+(2021); Chen et al.
+(2023); Tao et al.
+(2023); Zheng et al.
+(2023b) use probing techniques to measure forgetting in incremental learning.
+For example, Davari et al.
+(2022) uses linear probing to reveal representation drift due to parameter updates.
+Wu et al.
+(2021) conducts layer-wise probing on BERT, showing catastrophic forgetting in the top and middle layers.
+Chen et al.
+(2023) reveals a correlation between retaining past information and new task learning efficiency through linear probing on k-shot samples.
+Tao et al.
+(2023) illustrates BERT’s resilience to catastrophic forgetting even without buffer data.
+More recently, Zheng et al.
+(2023b) reveal that most previous work in NLP overlooks the data leakage issue and uses probing techniques to show that LLMs have superior performance on evaluated datasets even before incremental training.
+Additionally, they reveal that LLMs have strong anti-forgetting ability even under the sequential fine-tuning setting.
+Our work is inspired by Zheng et al.
+(2023b) and proposes a novel dataset to minimize the influence of data leakage issues, allowing for correct conclusions about the incremental learning ability of LLMs.
+A.2 Understanding Memorization
+Fewer studies Carlini et al.
+(2021); Tirumala et al.
+(2022); Biderman et al.
+(2024b); Boix-Adserà et al.
+(2023) explore memorization within LLMs.
+For instance, Carlini et al.
+(2021) discover that GPT-2 can memorize a small proportion of private information during pretraining, raising privacy concerns.
+Tirumala et al.
+(2022) show that larger LLMs memorize faster and have higher “forgetting baselines”.
+Biderman et al.
+(2024b) reveal the difficulty in predicting which training samples will be memorized by large language models.
+Boix-Adserà et al.
+(2023) find that transformers incrementally learn new knowledge, with trained and initial weights progressively increasing in rank.
+These studies explore memorization from both the perspective of sentences and model weights.
+In contrast, this paper studies the problem of memorization and forgetting at a more fine-grained level: the concept level.
+This approach addresses an underexplored research problem in the incremental learning community.
+A.3 Applications in NLP
+Many works Sun et al.
+(2020); Chuang et al.
+(2020); Liu and Huang (2023); Qiu et al.
+(2024); Zhang et al.
+(2023a); Shao et al.
+(2023a); Zhang et al.
+(2023b) focus on incremental learning for various NLP tasks, assuming catastrophic forgetting in pre-trained language models and designing techniques to mitigate it.
+These tasks include text classification Sun et al.
+(2020); Chuang et al.
+(2020), relation extraction Liu and Huang (2023), named entity recognition Qiu et al.
+(2024); Zheng et al.
+(2022); Zhang et al.
+(2023a), intent classification Shao et al.
+(2023a), and machine translation Zhang et al.
+(2023b).
+We refer readers to the survey Zheng et al.
+(2024) for more applications of incremental learning in NLP tasks.
+The proposed Concept-1K differs substantially from the aforementioned NLP tasks.
+It is more challenging and provides better explainability at the concept level.
+Appendix B Data Leakage in IL of LLMs
+The data leakage issue in NLP is often implicit.
+Unlike computer vision, where pretraining involves explicit category information, NLP pretraining is self-supervised and lacks clear categorical distinctions that can be easily compared between the pretraining corpus and downstream datasets.
+This makes it challenging to detect and address data leakage.
+Therefore, we urge future studies to exercise greater caution regarding data leakage in the IL of LLMs.
+B.1 Data Leakage in Classification Tasks
+The issue of data leakage in classification tasks for IL with LLMs has recently been highlighted by Zheng et al.
+(2023b).
+Their extensive study revisits over 20 IL methods across four key classification tasks: Text Classification, Intent Classification, Relation Extraction, and Named Entity Recognition.
+One core finding of Zheng et al.
+(2023b) is that LLMs, such as BERT and GPT-like models, exhibit high probing performance even before they are incrementally trained on specific downstream tasks.
+This high initial performance suggests that these models already possess substantial knowledge relevant to the classification tasks due to their extensive pre-training on diverse corpora.
+Consequently, when these LLMs are evaluated under IL settings, the incremental learning of new tasks may, in fact, be leveraging pre-existing knowledge rather than genuinely incremental learning.
+This phenomenon leads to misleading conclusions about the effectiveness of various IL methods Sun et al.
+(2020); Chuang et al.
+(2020); Liu and Huang (2023); Qiu et al.
+(2024); Zhang et al.
+(2023a); Shao et al.
+(2023a); Zhang et al.
+(2023b).
+The high probing performance before task-specific training indicates that the models are not learning incrementally as assumed but rather recalling previously acquired knowledge.
+Therefore, many IL studies in the context of classification tasks suffer from data leakage, as the benchmark tasks are not truly novel to the LLMs.
+Addressing this issue requires carefully designed benchmarks that ensure the novelty and exclusivity of the knowledge being tested, a challenge we aim to tackle with our Concept-1K dataset.
+B.2 Data Leakage in Generation Tasks
+Data leakage poses a significant challenge in IL for generation tasks, which are often more specific and diverse compared to classification tasks.
+For example, in the IL setting described by Scialom et al.
+(2022), the task sequence includes Text Simplification, Headline Generation with Constraints, Haiku Generation, Covid QA, Inquisitive Question Generation, Empathetic Dialogue Generation, Explanation Generation, and Twitter Stylometry.
+Despite the specificity and diversity of these tasks, data leakage remains a concern because LLMs are trained on extensive corpora from the internet and often undergo supervised finetuning on dialogue data OpenAI (2023).
+This pre-training on vast amounts of internet data means that LLMs might already possess significant knowledge relevant to these generation tasks.
+Moreover, the data leakage issue in generation tasks is implicit and easy to overlook.
+Unlike classification tasks, where techniques like probing Zheng et al.
+(2023b) can measure LLM performance before training, generation tasks lack such straightforward methods to assess initial model capabilities.
+This makes it challenging to ascertain how much new knowledge is genuinely being learned during IL versus what the model is simply recalling from its pre-trained knowledge base.
+Pioneering work investigating the IL ability of LLMs found that the T0 model Sanh et al.
+(2021) barely forgets when learning new tasks.
+This suggests that the T0 model may have already acquired the ability to solve multiple generation tasks with appropriate input prompts before explicit training on them.
+Detailed results are presented in Figure 2 of Sanh et al.
+(2021).
+Another study by Zhang et al.
+(2023c) defines generation tasks using different instructions, a paradigm they call continual instruction tuning.
+They train the T5 model Raffel et al.
+(2020) sequentially on 19 tasks, achieving 35.7% performance, while jointly training on these tasks yields 42.1%.
+The close gap between upper bound and lower bound performance indicates minimal forgetting and suggests potential data leakage.
+Detailed results and experimental settings are provided in Tables 1 and 2 of Zhang et al.
+(2023c).
+Although recent studies Module ; Guo et al.
+(2024); Huang et al.
+(2024); Yang et al.
+(2024); Peng et al.
+(2024); Ren et al.
+(2024) claim that forgetting is serious in continual instruction tuning, all of them utilize parameter-efficient finetuning techniques such as LoRA Hu et al.
+(2021) or prompt tuning Lester et al.
+(2021).
+As shown in our experiments in Section 3.3, the IL ability of LoRA and full finetuning differ substantially.
+LoRA significantly limits the ability to learn new concepts compared to full finetuning, leading to limited new knowledge acquisition and faster forgetting on training samples.
+Therefore, their IL settings may not accurately reflect the true IL ability of LLMs.
+Appendix C Comparison with Existing Incremental Learning Setting
+There are three popular IL settings which are widely adopted in the literature of computer vision: class-incremental learning (CIL), task-incremental learning (TIL), and continual pretraining (CPT).
+However, none of them are appropriate to evaluate the IL ability of LLMs.
+C.1 Class-Incremental Learning
+CIL is designed for classification tasks such as text classification, and its goal is to learn new classes incrementally.
+On the one hand, SOTA LLMs with billion-level parameters are overqualified for the above classification tasks with only dozens of categories.
+On the other hand, the pretraining corpus is likely to contain the knowledge required for the classification tasks (data leakage issue).
+As shown empirically by Zheng et al.
+(2023b), sequential training frozen LLMs with expanding classifiers yields comparable or even superior performance with SOTA IL methods.
+C.2 Task-Incremental Learning
+TIL aims to learn new tasks incrementally Sun et al.
+(2020); Qin and Joty (2022).
+Apart from the data leakage issue, the diversity of tasks and orders across research makes it difficult to readily and fairly compare IL algorithms.
+C.3 Continual Pretraining
+The last scenario CPT aims at continual pretraining models on the corpus from different domains.
+However, the evaluation relies on the performance of downstream tasks, where we can hardly identify what knowledge is learned or forgotten.
+C.4 Summary
+In this paper, we consider a novel IL scenario called Instance-level Incremental Learning (IIL).
+Unlike the IL scenario mentioned above, IIL regards each concept as an instance and is more practical and challenging for existing LLMs.
+Specifically, we are motivated by the human learning process and expect LLMs to learn new concepts incrementally without forgetting.
+For example, in Figure 1, humans can learn new concepts that are constantly emerging, such as “Metaverse” and “Quantum Computing”.
+After learning more concepts such as “Web3.0” and “Non-Fungible Token”, humans will not immediately forget the previously learned concepts such as “Metaverse”.
+Appendix D Introduction of Domains in Concept-1K
+The domains in Concept-1K are introduced as follows:
+D.1 Technology Domain
+This domain focuses on both cutting-edge and widely applied technologies.
+Cutting-edge technologies include artificial intelligence, blockchain, quantum computing, etc., while widely applied technologies encompass cloud computing, the Internet of Things, and more.
+D.2 Economic Domain
+This domain highlights economic trends and emerging economic concepts.
+Economic trends cover topics such as digital currency and globalization, whereas emerging concepts include quantitative computing, electronic wallets, peer-to-peer (P2P) networks, and others.
+D.3 Education Domain
+This domain emphasizes emerging educational technologies and concepts.
+Technologies such as remote learning and online courses, along with concepts like bilingual education, social education, and lifelong learning, are included.
+D.4 Environmental Domain
+This domain centers on global environmental issues and green technologies.
+Topics include climate change, environmental protection, and sustainable energy, as well as technologies like green roofs, shared bicycles, and solar panels.
+D.5 Cultural Domain
+This domain focuses on diversity and inclusion, and digital media and arts.
+Diversity and inclusion cover multiculturalism, gender equality, and social inclusion, while digital media and arts include digital art, social media trends, and online communities.
+D.6 Health and Medical Domain
+This domain is dedicated to emerging medical technologies and public safety.
+It covers CRISPR gene editing technology, the application of artificial intelligence in medical diagnosis, telemedicine services, wearable health monitoring devices, and concepts related to vaccine development, disease monitoring and prevention strategies, and promoting public health awareness.
+Appendix E Concept Selection Criterion
+In constructing Concept-1K, we carefully selected concepts based on the following criteria to ensure the relevance, novelty, and richness of the learning material:
+E.1 Length criterion
+The concept words should not exceed three words in length.
+This encourages the model to focus on significant and concise terms within the domain, facilitating efficient learning and ensuring that the concepts provide a rich source of information.
+Shorter concepts are easier to manage and help avoid potential confusion that may arise from overly complex or verbose terms.
+E.2 Timeliness criterion
+The chosen concept words should preferably be those that emerged after January 2022.
+This criterion ensures that the general pre-trained models have not yet encountered and learned these concepts and the associated knowledge.
+By selecting recent concepts, we aim to test the true incremental learning capabilities of LLMs, avoiding biases introduced by prior knowledge.
+E.3 Trend criterion
+We focus on concepts that are currently receiving widespread attention in academia, industry, and the media.
+This ensures that the selected concepts are not only relevant and contemporary but also significant and impactful in their respective fields.
+By choosing trending concepts, we can better gauge the models’ ability to learn and adapt to the latest advancements and discussions in various domains.
+Table 10: The statistics of the LLMs used in this paper. $\dagger$: Non-embedding parameters according to Biderman et al. (2023).
 Model Class Pretrained Weights Parameters Layers Hidden Dim Link
@@
 llama-2-13b-hf 13B 40 5120 Link
-Appendix B Experimental Settings
-Backbones.
+Appendix F Experimental Settings
+F.1 Backbones
 We use the Pythia suite Biderman et al.
-(2023) and the popular open-source models including LLaMa and Vicuna for experiments.
+(2023) and other popular open-source models, including LLaMa and Vicuna, for our experiments.
 Pythia is based on GPT-NeoX Black et al.
-(2022), which contains 8 model sizes and 154 pre-training checkpoints, enabling research in interpretability and learning dynamics.
-The statistics of the 9 PLMs used in this paper are summarised in Table 6.
+(2022) and includes 8 model sizes and 154 pre-training checkpoints, facilitating research in interpretability and learning dynamics.
+The statistics of the 9 LLMs used in this paper are summarized in Table 10.
 We download the pre-trained weights from Huggingface (Wolf et al., 2019).
-Datasets.
-We split the 1023 concepts in Concept-1K into 10 groups for IIL.
-Specifically, they sorted the concepts alphabetically and shuffled the order with random seed 1.
-The first group contains 105 concepts, while the others contain 102 concepts.
-Baseline Methods. The introduction of the baseline methods is as follows:
-  • SEQ: SEQ refers to sequential fine-tuning. It can be considered as the lower bound of incremental learning.
-  • REPLAY: Experience replay stores representative old samples and jointly optimizes old and new samples when learning new tasks. It is a practical and popular technique in incremental learning.
+F.2 Implementation Details
+We sort the concepts alphabetically and shuffle the order using random seed 1.
+The maximum input and output lengths are set to 32 and 10, respectively.
+The batch size is 32, and the learning rate for the LLMs is $1\times 10^{-5}$.
+We use the AdamW optimizer Loshchilov and Hutter (2018).
+For LLMs with more than 1B parameters, we use A800 GPUs, while smaller LLMs run on RTX3090 GPUs.
+Each experiment is repeated three times, and we report the average and standard deviations.
+Additionally, we search for the best hyper-parameters for each baseline method.
+For the experiment in Section 3.2, we use “gpt-3.5-turbo” and “gpt-4-turbo” for GPT-3.5 and GPT-4, respectively.
+Given the high cost of evaluating the entire Concept-1K dataset, we randomly sample 500 instances for both GPT-3.5 and GPT-4.
+The outputs and targets of GPT-3.5 and GPT-4 on these 500 instances are provided in the supplementary material.
+F.3 Input Prompt
+We use the following input prompt for training and testing Concept-1K:
+Question: {Question}
+Short Answer: {Answer} ,
+where {Question} and {Answer} represent the question and the target output, respectively. For the experiment in Section 3.2, the prompts are provided in Tables 11 and 12.
+Table 11: Prompt for In-Context Learning with 1 shot demonstration.
+{Question i} and {Answer i} represent the question and the target output of the $i$-th demonstration training sample respectively.
+{Test Question} represents the test question.
+I will provide some knowledge as follows:
+Question: {Question 1}
+Short Answer: {Answer 1}
+Please answer the following question according to the above knowledge:
+Question: {Test Question}
+Short Answer:
+Table 12: Prompt for In-Context Learning with 5 shot demonstrations.
+{Question i} and {Answer i} represent the question and the target output of the $i$-th demonstration training sample respectively.
+{Test Question} represents the test question.
+I will provide some knowledge as follows:
+Question: {Question 1}
+Short Answer: {Answer 1}
+Question: {Question 2}
+Short Answer: {Answer 2}
+Question: {Question 3}
+Short Answer: Answer 3
+Question: {Question 4}
+Short Answer: {Answer 4}
+Question: {Question 5}
+short Answer: {Answer 5}
+Please answer the following question according to the above knowledge:
+Question: {Test Question}
+Short Answer:
+Appendix G Introduction of Baseline Methods
+The introduction of the baseline methods is as follows:
+  • SEQ: Sequential fine-tuning (SEQ) is considered the lower bound of incremental learning.
+  • REPLAY: Experience replay stores representative old samples and jointly optimizes both old and new samples when learning new tasks.
+This is a practical and popular technique in incremental learning.
   • EWC Kirkpatrick et al.
-(2017): EWC is a regularization-based method.
-The weight of each parameter is determined by the diagonal of the Fisher information matrix.
-The weight of the regularization loss is searched in {5×$\times 10^{3}$, 1×$\times 10^{4}$, 5×$\times 10^{4}$, 1×$\times 10^{5}$, 1×$\times 10^{6}$, 1×$\times 10^{7}$}.
+(2017): Elastic Weight Consolidation (EWC) is a regularization-based method where the weight of each parameter is determined by the diagonal of the Fisher information matrix.
+The regularization loss weight is searched within {5×$\times 10^{3}$, 1×$\times 10^{4}$, 5×$\times 10^{4}$, 1×$\times 10^{5}$, 1×$\times 10^{6}$, 1×$\times 10^{7}$}.
   • LAMOL Sun et al.
-(2020): LAMOL trains PLMs with question-answering and generative objectives.
-It generates pseudo-samples before learning each new task for data replay.
-The weight of the generation loss is $\lambda=0.25$.
-The proportion of pseudo samples is $\gamma=0.20$.
-There are two variants of LAMOL, including LAMOL_t and LAMOL_g.
-The difference is whether a task-specific token is used for generation.
+(2020): LAMOL trains LLMs with question-answering and generative objectives, generating pseudo-samples before learning each new task for data replay.
+The generation loss weight is $\lambda=0.25$, and the proportion of pseudo-samples is $\gamma=0.20$.
+There are two variants: LAMOL_t and LAMOL_g, differing by whether a task-specific token is used for generation.
   • L2KD Chuang et al.
-(2020): L2KD adds a knowledge distillation target based on LAMOL.
-The teacher model is trained from scratch.
-We implemented the word-level variant since it performs best on text classification tasks.
+(2020): L2KD adds a knowledge distillation target based on LAMOL, with the teacher model trained from scratch.
+We implemented the word-level variant as it performs best on text classification tasks.
   • LAMOL_KD Zheng et al.
-(2023): LAMOL_KD utilizes knowledge distillation based on LAMOL_t.
+(2023b): LAMOL_KD utilizes knowledge distillation based on LAMOL_t.
 Unlike L2KD, the teacher model in LAMOL_KD is trained on all previous tasks.
-The new data are used to learn the LAMOL objectives, and pseudo data are used for word-level knowledge distillation as a regularization term.
-  • PCLL Zhao et al. (2022): PCLL combines the idea of variational autoencoders and word-level knowledge distillation with the objective of LAMOL.
-  • LFPT5 Qin and Joty (2022): LFPT5 only learns soft prompts for each new task. The training objective is the same as LAMOL. The number of soft prompt tokens is 10.
+New data are used to learn the LAMOL objectives, and pseudo data are used for word-level knowledge distillation as a regularization term.
+  • PCLL Zhao et al. (2022): PCLL combines the concepts of variational autoencoders and word-level knowledge distillation with the objectives of LAMOL.
+  • LFPT5 Qin and Joty (2022): LFPT5 learns only soft prompts for each new task. The training objective is the same as LAMOL. The number of soft prompt tokens is 10.
   • LoRA Hu et al.
-(2021): LoRA only trains a small proportion of parameters of PLMs.
+(2021): LoRA trains a small proportion of parameters of LLMs.
 We set the rank $r=8$ and the scaling parameter $\alpha=16$.
-We use the implementation of LoRA from the PEFT library Mangrulkar et al.
+We use the LoRA implementation from the PEFT library Mangrulkar et al.
 (2022).
-We note that some IL methods are not compared since they are not applicable in the IIL scenario.
+Some IL methods are not compared as they are not applicable in the IIL scenario.
 For example, Progressive Prompt Razdaibiedina et al.
-(2023) requires task ID during both training and inference stages.
+(2023) requires task IDs during both training and inference stages.
 VAG Shao et al.
-(2023) requires storing the vocabulary of class labels and do not apply to the generation tasks without class labels.
-Besides, prompt-based IL methods such as L2P Wang et al.
-(2022) also do not apply to generation tasks.
-Implementation Details.
-The maximum input and output lengths are set as 32 and 10, respectively.
-The batch size is 32.
-The learning rate of PLMs is $1\times 10^{-5}$.
-We use AdamW Loshchilov and Hutter (2018) optimizer.
-We use A800 GPUs for our experiments.
-Each experiment is repeated three times, and the average and standard deviations are reported.
-We search for the best hyper-parameters for each baseline method.
-Figure 3: The accuracy of different concepts in the first task after sequential finetuning Pythia-2.8B on the 10 tasks of Concept-1K.
-Appendix C Additional Experimental Results and Analysis
-C.1 The Role of Model Scale, Pretraining, and Buffer Size
-We provide the detailed result of Figure 2 in Table 7, 8, 9, 10, 11, 12, 13, and 14.
-C.2 The performance of SOTA methods
-The backbone model is Pythia-410M.
-We provide the step-wise average accuracy on training and test set in Figure 4 (a) and (b).
-The results demonstrate a large gap between SOTA IL methods and data replay.
-We also provided a comparison between SOTA methods in terms of memorization accuracy, generalization accuracy, memorization forgetting, and generalization forgetting in Figure 4 (c), (d), (e), and (f).
-In addition, from Figure 4 (g), we find that the training loss of the prompt-tuning-based method LFPT5 can not decrease to a small value.
-It indicates that only finetuning prompts are not practical for learning new knowledge, which is also consistent with the result of LoRA.
-C.3 In-Context Learning
-Apart from the analysis in the main paper, Table 5 shows that the smallest PLM (Pythia-70M) achieve high accuracy under the “1 shot” and “Same Know.” settings.
-The reason is that small PLMs simply copy the output in the demonstration instance as the final output.
-Under the “5 shot” and “Same Know.” settings, the accuracy of Pythia-70M is only 15.02%.
-We use the model “gpt-3.5-turbo” and “gpt-4-turbo” for GPT3.5 and GPT4, respectively.
-Due to the high cost of evaluating on the whole concept-1K, we randomly sample 500 instances for GPT3.5 and GPT4.
-The output and the target of GPT3.5 and GPT4 on the 500 instances are provided in the supplementary material.
-Table 5 shows that GPT4 achieves 86.20% under the “5 shot” and “Same Know.” settings.
-It indicates that the training and the test instance of Concept-1K share the same knowledge.
-C.4 Concept-Level Accuracy on Concept-1K
-We analyze the memorization accuracy of the concepts in the first task after IIL.
-The backbone model is Pythia-2.8B, and it is sequential finetuning on Concept-1K.
-From Figure 3, we have the following observation:
-  • The concepts such as “Cultural Competence”, “Augmented Reality Displays”, “Geopolitical Risks” that achieve high accuracy are broad, interdisciplinary topics that are likely rich in diverse data sources and content.
-These subjects span various fields and are likely to have clear definitions, significant public interest, and extensive coverage in the pretraining datasets, facilitating high memorization accuracy.
-  • The lowest accuracy is observed in highly technical, cutting-edge, or niche fields such as “Spatial Audio”, “Biomaterials”, “Sustainable Tech”, and “Space Tourism”.
-These topics are likely challenging due to rapidly evolving information, specialized knowledge requirements, and potentially limited representation in available datasets.
-In summary, the tendency from high to low accuracy reflects the transition from broad, interdisciplinary topics with abundant and diverse data to more specialized, technical, or emerging fields with potentially limited data exposure.
-This suggests that the PLMs’ memory accuracy is influenced by the richness and diversity of information available in its training data and the clarity and stability of the concepts themselves.
-Additionally, despite their significance, emerging technologies and areas of innovation might be harder for PLMs to learn accurately due to the rapid evolution of knowledge and possibly less structured data.
+(2023b) requires storing the vocabulary of class labels and does not apply to generation tasks without class labels.
+Additionally, prompt-based IL methods such as L2P Wang et al.
+(2022b) are not suitable for generation tasks.
+(a) After Task 1
+(b) After Task 2
+(c) After Task 4
+(d) After Task 6
+(e) After Task 8
+(f) After Task 10
+Figure 7: The visualization of the memorized knowledge related to the “Brain-Computer Interface” in IL.
+The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding training samples are answered correctly.
+(a) After Task 1
+(b) After Task 2
+(c) After Task 4
+(d) After Task 6
+(e) After Task 8
+(f) After Task 10
+Figure 8: The visualization of the generalized knowledge related to the “Brain-Computer Interface” in IL.
+The center node represents a concept, while the linked and unlinked edges indicate whether the corresponding test samples are answered correctly.
+Appendix H Visualization of One Concept
+We visualize the memorized and generalized knowledge of the concept “Brain-Computer Interface” in Figures 7 and 8.
+In each graph, the center node represents the concept “Brain-Computer Interface”.
+The linked and unlinked edges indicate whether the corresponding training or test samples are answered correctly.
+For example, in Figure 7a, the edge between “Brain-Computer Interface” and “Signal Processing” signifies that the LLM correctly outputs the target answer “Signal Processing” when the question is the training sample “What subevent occurs in a Brain-Computer Interface?”.
+In Figure 7b, the edge between “Brain-Computer Interface” and “Signal Processing” is missing, indicating that the LLM fails to provide the correct target answer.
+Figure 7 shows that even when LLMs can memorize all the knowledge, they tend to first forget more complex knowledge, such as “(Brain-Computer Interface, UsedFor, Controlling Computers With Thought)”.
+Conversely, some common knowledge, such as “(Brain-Computer Interface, Requires, Brain Signals)”, “(Brain-Computer Interface, Has Property, Innovative)”, and “(Brain-Computer Interface, Motivated By Goal, Accessibility)”, remains robust and is not forgotten after learning 10 tasks.
+This indicates that certain knowledge is easier to memorize, generalize, and retain.
+Further exploration at the concept-level knowledge in IL is left for future work.
+We also encourage future studies to utilize the provided Concept-1K dataset for a fine-grained analysis of the memorization and generalization dynamics in IL.
+Appendix I Additional Experimental Results
 Table 13: The memorization accuracy when different model scales and buffer size are selected. The pretraining step is 143000 (final version). Figure 4 (a) summarises this figure’s content.
@@
 What aspect of health does Nutritional Genomics have an impact on?
-Table 16: (Continual) Examples of Concept-1K. Each triplet corresponds to a training instance and a test instance.
-Domain Concept Triplet Training and Test Input Target Output
-Environment Groundwater Recharge (Groundwater Recharge, IsA, HydrologicalProcess) What is Groundwater Recharge classified as? hydrological process
-What kind of process is Groundwater Recharge?
-(Groundwater Recharge, UsedFor, AquiferSustainability) What is Groundwater Recharge used for? aquifer sustainability
-What purpose does Groundwater Recharge serve in relation to aquifers?
-(Groundwater Recharge, Requires, PermeableSurfaces) What does Groundwater Recharge require? permeable surfaces
-What are essential for the process of Groundwater Recharge?
-(Groundwater Recharge, ResultsIn, WaterTableRise) What is a result of Groundwater Recharge? water table rise
-What does Groundwater Recharge lead to regarding water tables?
-(Groundwater Recharge, MotivatedByGoal, DroughtMitigation) What goal motivates groundwater recharge? drought mitigation
-Why is groundwater recharge important?
-Sea Level Rise (Sea Level Rise, CausedBy, GlobalWarming) What causes Sea Level Rise? global warming
-What is the primary factor leading to Sea Level Rise?
-(Sea Level Rise, AnalyzedBy, Climatologists) Who analyzes Sea Level Rise? climatologists
-What group of professionals study Sea Level Rise?
-(Sea Level Rise, ResultsIn, HabitatLoss) What does Sea Level Rise result in? habitat loss
-What is a significant impact of Sea Level Rise on natural habitats?
-(Sea Level Rise, MeasuredBy, TideGauges) How is sea level rise measured? tide gauges
-What instrument is used to measure sea level rise?
-(Sea Level Rise, AddressedBy, EmissionReductions) How is Sea Level Rise addressed? emission reductions
-What strategy addresses Sea Level Rise?
 Figure 10: The histogram of the relations in Concept-1K. The top-50 relations are shown.